### v1

# Limit-case admissibility for positive infinite-dimensional systems

Abstract
In the context of positive infinite-dimensional linear systems, we systematically study $L^{p}$-admissible control and observation operators with respect to the limit-cases $p=\infty$ and $p=1$, respectively. This requires an in-depth understanding of the order structure on the extrapolation space $X_{-1}$, which we provide. These properties of $X_{-1}$ also enable us to discuss when zero-class admissibility is automatic. While those limit-cases are the weakest form of admissibility on the $L^{p}$-scale, it is remarkable that they sometimes follow from order theoretic and geometric assumptions. Our assumptions on the geometries of the involved spaces are minimal.
keywords:
Positive system; positive semigroup; Banach lattice; ordered Banach space; extrapolation space; admissible control operator; admissible observation operator; infinite-dimensional linear systems
{MSCcodes}
93C25, 93C05, 93C28, 47D06, 46B42, 47B65
1 Introduction
In this paper, we study the boundedness of linear operators
$\Phi_{\tau}:L^{p}([0,\tau];U)\to X,\qquad u\mapsto x(\tau),\quad\tau>0$
for some $p\in[1,\infty]$, arising in boundary control systems [24, 46, 50] of the form
$\dot{x}(t)={}$ $\mathfrak{A}x(t),\quad t>0,\quad x(0)=0,$
$\mathfrak{B}x(t)={}$ $u(t).$
Here $\mathfrak{A}:\operatorname{dom}\mathfrak{A}\subset X\to X$ and $\mathfrak{B}:\operatorname{dom}\mathfrak{A}\to U$ are linear operators acting on Banach spaces $X$ and $U$. Under the assumptions that the restriction $A=\mathfrak{A}|_{\ker\mathfrak{B}}$ generates a $C_{0}$-semigroup and that $\mathfrak{B}$ has a bounded right-inverse, the above-mentioned boundedness can be rephrased in terms of admissible operators, going back to Weiss [55, 54]. We also refer to [16], where the equivalent viewpoint used in this introduction was first taken. Indeed, the boundedness of $\Phi_{\tau}$ is equivalent to the property that for some $\lambda$ in the resolvent set $\rho(A)$ of $A$, the operator
$\widetilde{\Phi}_{\lambda,\tau}:L^{p}([0,\tau];U)\to X,\qquad u\mapsto\int_{0}^{\tau}T(\tau-s)B_{\lambda}u(s)\mathrm{ds}$
has range in $\operatorname{dom}A$ for some (hence all) $\tau>0$, where $B_{\lambda}=\left(\mathfrak{B}|_{\ker(\mathfrak{A}-\lambda)}\right)^{-1}$ is a well-defined bounded operator from $U$ to $X$, see e.g. [24] or [16, Remark 2.7]. In that case, $\Phi_{\tau}u=x(\tau)=(A-\lambda)\widetilde{\Phi}_{\lambda,\tau}u$ for any $\lambda\in\rho(A)$.
Admissible operators are an indispensable tool in the study of infinite-dimensional systems and particularly in the context of well-posed systems [30, 51, 49]. We point out that in the literature the definition of admissible (control) operators is usually given for systems in state-space form,
(1.1) $\dot{x}(t)=Ax(t)+Bu(t),\quad x(0)=0.$
This class includes – via the choice $B=(\lambda-A_{-1})B_{\lambda}$ with $A_{-1}$ being the extension of $A$ to the extrapolation space $X_{-1}$ corresponding to the semigroup – boundary control systems if $B$ is only required to map $U$ to $X_{-1}$, see [46] or [48, Section 2]. Admissibility for $p=2$ confines a rich theory to this date, if $X$ is a Hilbert space, see, for instance, [29, 49, 50]. While several results for $p\in(1,\infty)\setminus\{2\}$ exist [26, 25, 31, 54], the case $p=\infty$ has been studied systematically only recently [28, 33, 56]. We point out that the methods required to study admissibility strongly depend on the setting; such as whether or not $X$ is a Hilbert space, or whether the semigroup is analytic or can be extended to a group.
In this note we focus on the case $p=\infty$ and add another additional structure to the setting: positivity of the semigroup and the operator $B$ (or $B_{\lambda}$) – in the sense of ordered Banach spaces. In fact, the positivity of $B$ can be characterised by the positivity of the operators $B_{\lambda}$, see Subsection 4.1. While positivity is a well-studied concept for operator semigroups, its relation to admissibility, or more general infinite-dimensional systems theory, is less understood. In [16, Proposition 4.3], it is shown that positivity of $\Phi_{\tau}$ and $B_{\lambda}$ for sufficiently large $\lambda$ are equivalent provided that admissibility is assumed. Controllability of positive systems is studied in [16, 18]. Admissibility criteria for positive $B$ are given in [56, Chapter 4] and [19, Theorem 2.1]. The first reference focuses on $L^{\infty}$- and $C$-admissibility by imposing assumptions mainly on the space $U$. In the second reference, a strong assumption on the semigroup is made which implies $L^{1}$-admissibility for all positive $B$’s. In Appendix B, we provide a proof that under a reasonable compactness assumption, this condition can only be satisfied if $X$ is an $L^{1}$-space (Corollary B.3). Our aim – in contrast – is to list assumptions on the state space such that admissibility follows from the assumed positivity in a rather automatic fashion. First results in this direction were derived by Wintermayr [56]. It should be noted that even the notion of positivity for control operators $B$ stemming from (1.1) is nontrivial as positivity in the space $X_{-1}$ needs to be defined suitably [10]. Motivated by this, we devote Section 2 to a refined study of the order structure of $X_{-1}$, which is of interest in its own right. Under mild assumptions, the positive cone $X_{+}$ turns out to be a face in the cone $X_{-1,+}$. Moreover, we show that if the semigroup under consideration satisfies sufficiently strong ultracontractivity assumptions, then the cone $X_{-1,+}$ is contained in an interpolation space between $X$ and $X_{-1}$ – which immediately yields admissibility (Section 4.4).
In addition, we also study the formally dual notion of $L^{1}$-admissible observation operator $C:\operatorname{dom}\to Y$, with $Y$ being another Banach space, which means that the mapping
$\Psi_{\tau}:\operatorname{dom}A\to L^{1}([0,\tau];Y),x_{0}\mapsto CT(\cdot)x_{0}$
extends to a bounded linear operator on $X$. We present two types of results: in the first type, we assume conditions on the order structure of the spaces $X$ and $Y$, as well as positivity of the semigroup and observation operator. Another result, which is independent of any positivity assumptions and thus of interest in its own right, is Theorem 3.13, stating that zero-class $L^{1}$-admissibility is automatic from $L^{1}$-admissibility if $X$ is reflexive and $Y$ an $\mathrm{AL}$-space. The case of $L^{\infty}$-admissible control operators is dealt with in Section 4, which also generalises several results from Wintermayr [56]. The derived results also relate to input-to-state stability of infinite-dimensional systems, a notion that has seen tremendous research interest within the past decade; see [39] for an overview.
We point out that in the limiting cases of the Hölder conjugates $p=1$ and $p=\infty$ the duality between admissible control and observation operators is subtle and does not allow for a one-to-one translation between admissibility for observation operators and control operators.
In the last section, Section 5, we apply our findings to certain well-known perturbation results. Certain assumptions of Sections 3 and 4 are elaborated on in Appendix A. In the rest of the introduction, we recall basic concepts and fix our notations.
Banach spaces and operators
The closed unit ball of a Banach space $X$ is denoted by $B_{X}$. The space of bounded linear operators between two Banach spaces $X$ and $Y$ will be denoted by $\mathcal{L}(X,Y)$ and for any linear operator $T$, by $\operatorname{dom}T,\ker T$, and $\operatorname{Rg}T$ we mean the domain and the range of $T$ respectively. Moreover, for a closed operator $A$ on $X$, we write $A^{\prime}$ for the dual operator of $A$ that acts on the dual Banach space $X^{\prime}$. The resolvent set and spectrum of $A$ are as usual denoted by $\rho(A)$ and $\sigma(A)$ respectively. Moreover, by $R(\lambda,A)$, we mean the resolvent of $A$ at a point $\lambda\in\rho(A)$ and the spectral bound of $A$ is given by
$s(A):=\sup\{\operatorname{Re}\lambda:\lambda\in\sigma(A)\}.$
Operator semigroups
We assume the reader is familiar with the theory of $C_{0}$-semigroups, for which we refer to the monograph [17]. The interpolation and extrapolation spaces associated with a $C_{0}$-semigroup play an important role in the context of admissibility. Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$ whose generator is denoted by $A$ and fix $\lambda\in\rho(A)$. The corresponding interpolation space
$X_{1}:=(\operatorname{dom}A,\left\lVert\mathord{\,\cdot\,}\right\rVert_{1})\quad\text{where}\quad\left\lVert\mathord{\,\cdot\,}\right\rVert_{1}:=\left\lVert\mathord{\,\cdot\,}\right\rVert+\left\lVert A\mathord{\,\cdot\,}\right\rVert$
and extrapolation space – defined as the completion
$X_{-1}:=(X,\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1})^{\sim}\quad\text{where}\quad\left\lVert\mathord{\,\cdot\,}\right\rVert_{-1}:=\left\lVert R(\lambda,A)\mathord{\,\cdot\,}\right\rVert$
with $R(\lambda,A):=(\lambda-A)^{-1}$, are both Banach spaces. Moreover, for different choices of $\lambda$, the norms on $X_{-1}$ are equivalent. Further, $(T(t))_{t\geq 0}$ extends uniquely to a $C_{0}$-semigroup on $X_{-1}$, denoted by $(T_{-1}(t))_{t\geq 0}$. The generator $A_{-1}$ of $(T_{-1}(t))_{t\geq 0}$ has domain $\operatorname{dom}A_{-1}=X$ and is the unique extension of $A$ to an isometry from $X$ onto $X_{-1}$. An important property that we use throughout without citing is the isomorphisms $(X_{1})^{\prime}=(X^{\prime})_{-1}$ and $(X_{-1})^{\prime}=(X^{\prime})_{1}$, where $(X^{\prime})_{-1}$ and $(X^{\prime})_{1}$ refer to the dual semigroup on $X^{\prime}$; see [53, Corollary 3.1.17].
Ordered Banach spaces and Banach lattices
In concrete examples, usually the state space of a system is a function space such as $L^{p}$ and hence a Banach lattice. Yet, in a system theoretic context, natural candidates for input or output spaces are often spaces of differentiable functions such as, for instance, Sobolev spaces on the boundary; these are ordered Banach spaces but not Banach lattices. Moreover, it was shown in [10, Example 5.1] that, even if one starts with a Banach lattice $X$, the canonical order on the extrapolation space $X_{-1}$ may not render $X_{-1}$ a Banach lattice. For these reasons, we find it more natural to set up the entire theory within the framework of ordered Banach spaces.
An ordered Banach space $X$ is actually an ordered pair $(X,X_{+})$, where $X$ is a Banach space and $X_{+}\subseteq X$ is a closed convex cone satisfying $X_{+}\cap(-X_{+})=\{0\}$. The subset $X_{+}$ is called the positive cone of $X$ because it induces a natural order on $X$$x\leq y$ if and only $y-x\in X_{+}$. Whenever referring to the cone of an ordered Banach space, we mean the positive cone. The set $X_{+}-X_{+}$ is always a vector subspace of $X$ and coincides with the linear span of $X_{+}$. The cone $X_{+}$ is called generating if $X=X_{+}-X_{+}$ and normal if there exists $M\geq 1$ such that for each $x,y\in X_{+}$ the inequality $x\leq y$ implies $\left\lVert x\right\rVert\leq M\left\lVert y\right\rVert$.
If $X$ is an ordered Banach space one can always endow the span $X_{+}-X_{+}$ with the norm
(1.2) $\left\lVert x\right\rVert_{X_{+}-X_{+}}:=\inf\{\left\lVert y\right\rVert+\left\lVert z\right\rVert:\,y,z\in X_{+}\text{ and }x=y-z\};$
which is a complete norm on $X_{+}-X_{+}$ stronger than the norm induced by $X$ and both norms coincide on $X_{+}$ [7, Lemma 2.2]. In particular, the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{X_{+}-X_{+}}$ turns $X_{+}-X_{+}$ into an ordered Banach space and if the cone $X_{+}$ is normal with respect to the norm in $X$, then it is also normal with respect to this new norm.
A subset $S$ of $X$ is said to be order-bounded if there exists $x,z\in X$ such that $S$ is contained in the so-called order interval $[x,z]:=\{y\in X:\,x\leq y\leq z\}$. A non-empty set $C\subseteq X_{+}$ is called a face of $X_{+}$ if $\alpha C+\beta C\subseteq C$ for all $\alpha,\beta\geq 0$ and $[0,x]\subseteq C$ for all $x\in C$. A subspace $V$ of $X$ is said to be majorizing in $X$ if for each $x\in X$, there exists $v\in V$ such that $x\leq v$.
An element $e\in X_{+}$ is called a unit if for each $x\in X$, there exists $c>0$ such that $x\leq ce$. Actually, $e\in X_{+}$ is a unit if and only if $e$ is an interior point of $X_{+}$, see for instance [22, Proposition 2.11] for more details. Therefore, whenever $X$ has a unit, we can – and will – choose an equivalent norm on $X$ such that the closed unit ball of $X$ is upwards directed, i.e., for each $x_{1},x_{2}\in B_{X}$ there exists $x\in B_{X}$ such that $x_{1},x_{2}\leq x$. If $e$ is a positive element of an ordered Banach space $X$, then the principal ideal generated by $e$ is defined as
(1.3) $X_{e}:=\bigcup_{\lambda>0}[-\lambda e,\lambda e].$
In addition, if the cone of $X$ is normal, then by [4, Theorem 2.60], $X_{e}$ becomes an ordered Banach space (with unit $e$) when equipped with the norm
$\left\lVert x\right\rVert_{e}:=\inf\{\lambda>0:x\in[-\lambda e,\lambda e]\}.$
{tblr}
[…]
Table 1: Important properties of ordered Banach spaces
Banach lattices form a special class of ordered Banach spaces. An ordered Banach space $X$ is called a Banach lattice if any two elements have a supremum (equivalently: infimum) and
$\left\lvert x\right\rvert\leq\left\lvert y\right\rvert\Rightarrow\left\lVert x\right\rVert\leq\left\lVert y\right\rVert;$
where $\left\lvert x\right\rvert=\sup\{x,-x\}$. The cone of a Banach lattice is always generating and normal. The most common examples of Banach lattices are the $L^{p}$ spaces for $p\in[1,\infty]$ and the spaces of continuous functions with the supremum norm. The theory of ordered Banach spaces and Banach lattices is classical and can be found in, for instance, [47, 38, 3, 1].
Many ordered Banach spaces $X$ have additional geometric properties and if $X$ is even a Banach lattice, those properties often have special names. Some of those properties are very useful in our results. To help the reader to keep an overview, we summarise those properties – along with their name in the case that $X$ is a Banach lattice and along with some common examples – in Table 1.
Let $T:X\to Y$ be a linear map between ordered Banach spaces $X$ and $Y$. We say that $T$ is positive and write $T\geq 0$ if $TX_{+}\subseteq Y_{+}$ and denote $\mathcal{L}(X,Y)_{+}$ by the positive, bounded linear operators between $X$ and $Y$. If $X_{+}$ is generating, then we automatically have $T\in\mathcal{L}(X,Y)$ [1, Corollary 6]. A functional $\varphi\in X^{\prime}$ is called strictly positive if its kernel contains no positive non-zero element. A linear operator $J:X\to Y$ between ordered Banach spaces is called bipositive if for each $x\in X$ we have $Tx\geq 0$ if and only if $x\geq 0$. It is easy to check that if $J$ is bipositive, then it is injective. Intuitively, if we have a bipositive linear map between ordered Banach spaces $X$ and $Y$ this means that we can consider $X$ as a subspace of $Y$ that is endowed with the order structure inherited from $Y$. Therefore, we shall sometimes say, by abuse of terminology, that $X_{+}$ is a face of $Y_{+}$ to mean that $J(X_{+})$ is a face of $Y_{+}$.
A $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$ is called positive if each operator $T(t)$ is positive. For the theory of positive $C_{0}$-semigroups, we refer the reader to [41, 11, 12].
Complexifications
The theories of ordered Banach spaces and Banach lattices are theories over the real field. However, whenever one intends to use spectral theoretic arguments it is natural to use complex scalars. To this end one can use complexifications of real Banach spaces as they are, for instance, described in [40], [21, Appendix C], and – specifically for the case of Banach lattices – [47, Section II.11]. The details do not cause any problems in our setting, so we just remark here that, whenever we discuss spectral properties of a linear operator $A$ between two real Banach spaces $X$ and $Y$, we shall tacitly always mean the property of the extension of $A$ to complexifications of $X$ and $Y$.
2 Order properties of the extrapolation space $X_{-1}$
To utilise any order assumptions on the control operator $B\in\mathcal{L}(U,X_{-1})$ in (4.1), it is crucial to understand the order structure of the space $X_{-1}$. In Section 2.1, we first show how positive bijections between order Banach spaces preserve important order properties. Subsequently, in Section 2.2, we introduce an order on $X_{-1}$ and apply the results of Section 2.1 to show how various order properties of $X$ are carried over to $X_{-1}$.
2.1 Transferring properties between ordered Banach spaces
The general setting of this subsection is two ordered Banach spaces $Z$ and $X$ that are related via a positive bijection $T\in\mathcal{L}(Z,X)$. We begin by proving some properties of bipositive operators between ordered Banach spaces.
Proposition 2.1.
Let $X$ and $Z$ be ordered Banach spaces and let $J\in\mathcal{L}(X,Z)$ be bipositive. The following assertions hold.
[…]
    The map $J$ is injective.
[…]
    If $J(X_{+})$ is dense in $Z_{+}$, then $J^{\prime}$ is bipositive.
[…]
    If $Z_{+}$ is generating in $Z$ and $J(X)$ is majorizing in $Z$, then $J^{\prime}(Z^{\prime}_{+})=X^{\prime}_{+}$
Proof 2.2.
(a) Let $x\in X$ and $Jx=0$. Then $Jx\geq 0$ and $J(-x)\geq 0$, so the bipositivity of $J$ implies that $x\geq 0$ and $-x\geq 0$. Hence $x=0$ since $X_{+}\cap-X_{+}=\{0\}$.
(b) Since $J$ is positive, so is $J^{\prime}$. Now let $z^{\prime}\in Z^{\prime}$ and assume that $J^{\prime}z^{\prime}\geq 0$. For every $x\in X_{+}$ this implies that $\langle z^{\prime},Jx\rangle\geq 0$. Since $J(X_{+})$ is by assumption dense in $Z_{+}$, we conclude that $\langle z^{\prime},z\rangle\geq 0$ for all $z\in Z_{+}$, so $z^{\prime}\in Z^{\prime}_{+}$.
(c) Since $J^{\prime}$ is positive, we only need to show the inclusion $J^{\prime}(Z^{\prime}_{+})\supseteq X^{\prime}_{+}$, so let $x^{\prime}\in X^{\prime}_{+}$. The linear mapping $\varphi:J(X)\to\mathbb{R}$, $J(x)\mapsto\langle x^{\prime},x\rangle$ is well-defined and positive by the injectivity – known from (a) – and bipositivity of $J$. Since $J(X)$ is by assumption majorizing in $Z$, the Kantorovich extension theorem [4, Theorem 1.36] implies $\varphi$ can be extended to a positive linear functional $z^{\prime}:Z\to\mathbb{R}$. As $Z_{+}$ is generating, it follows that $z^{\prime}$ is automatically continuous [1, Corollary 6], i.e., $z^{\prime}\in Z^{\prime}_{+}$. Moreover, the equality $\langle z^{\prime},Jx\rangle=\varphi(Jx)=\langle x^{\prime},x\rangle$ for each $x\in X$ implies that $x^{\prime}=J^{\prime}z^{\prime}\in J^{\prime}(Z^{\prime}_{+})$.
We note in passing that the proof of (b) only used the positivity rather than the bipositivity of $J$. The following classical example is simple but illustrates Proposition 2.1 quite well:
Example 2.3 (Dual cone of $\mathrm{C}^{k}(\overline{\Omega})$).
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain, let $k\in\mathbb{N}$, and let $\mathrm{C}^{k}(\overline{\Omega})$ denote the space of all $k$-times continuously differentiable functions on $\Omega$ whose derivative up to order $k$ extend continuously to $\overline{\Omega}$. This is an ordered Banach space when endowed with the pointwise order and the usual norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\mathrm{C}^{k}}$-norm. We show that its dual cone is (more precisely, can be identified with) the set of all positive finite Borel measures on $\overline{\Omega}$.
Indeed, the canonical embedding $J:\mathrm{C}^{k}(\overline{\Omega})\to\mathrm{C}(\overline{\Omega})$ is bipositive and has majorizing range because it contains the constant function $\operatorname{{\mathbbm{1}}}$. Moreover, employing the density of $\mathrm{C}^{k}(\overline{\Omega})$ in $\mathrm{C}(\overline{\Omega})$ (which is true by the Stone-Weierstraß approximation theorem) and by using shifts by small multiples of the constant function $\operatorname{{\mathbbm{1}}}$, we even have that the cone $\mathrm{C}^{k}(\overline{\Omega})_{+}$ is dense in the cone $\mathrm{C}(\overline{\Omega})_{+}$.
So the dual operator $J^{\prime}$ is bipositive and injective according to Proposition 2.1. Hence, $J^{\prime}$ is a bipositive embedding of the dual space $\mathrm{C}(\overline{\Omega})^{\prime}$ (which can be identified with the space of finite Borel measures on $\overline{\Omega}$) into the dual space of $\mathrm{C}^{k}(\overline{\Omega})$. Hence, Proposition 2.1(c) shows that the cone $\mathrm{C}^{k}(\overline{\Omega})$ consists of (to be precise, can be identified via $J^{\prime}$ with) the set of all positive finite Borel measures on $\overline{\Omega}$. In other words, there are not more positive functionals on $\mathrm{C}^{k}(\overline{\Omega})$ than on $\mathrm{C}(\overline{\Omega})$ since the former space is majorizing in the latter.
In our next result, we show that the first two properties in Table 1 are preserved by the inverse of a positive bounded bijection.
Theorem 2.4.
Let $Z$ and $X$ be ordered Banach spaces and assume that there exists a bijection $T\in\mathcal{L}(Z,X)_{+}$.
[…]
    If every positive, increasing, and norm-bounded net in $X$ is norm-convergent, then the same is true in $Z$.
[…]
    If the cone of $X$ is the face of its bidual cone, then the same is true for the cone of $Z$.
[…]
    If the cone of $X$ is normal, then so is the cone of $Z$.
Proof 2.5.
(a) Suppose that every increasing norm-bounded net in $X_{+}$ is norm-convergent and let $(z_{\alpha})$ be an increasing norm-bounded net in $Z_{+}$. Then the net $(Tz_{\alpha})$ is increasing and norm-bounded in $X_{+}$, and hence it is norm-convergent, i.e., there exists $x\in X$ such that $Tz_{\alpha}\to x$ in $X$. As $T$ is a continuous linear bijection, it follows by the bounded inverse theorem that $z_{\alpha}\to T^{-1}x$ in $Z$, as desired.
(b) Let $k:X\to X^{\prime\prime}$ and $\widetilde{k}:Z\to Z^{\prime\prime}$ be the canonical embeddings. Let $z\in Z_{+}$ and $z^{\prime\prime}\in Z^{\prime\prime}_{+}$ such that $0\leq z^{\prime\prime}\leq\widetilde{k}(z)$. We need to show that $z^{\prime\prime}\in\widetilde{k}(Z)$. Since $T$ is positive, so is the double dual $T^{\prime\prime}:Z^{\prime\prime}\to X^{\prime\prime}$. Therefore, $0\leq T^{\prime\prime}z^{\prime\prime}\leq T^{\prime\prime}\widetilde{k}(z)=k(Tz)$. We may now infer from the assumption, $k(X_{+})$ is a face of $X^{\prime\prime}_{+}$, that $T^{\prime\prime}z^{\prime\prime}\in k(X)$. It follows that $z^{\prime\prime}\in(T^{\prime\prime})^{-1}k(X)=\widetilde{k}(T^{-1}X)=\widetilde{k}(Z)$; here we’ve implicitly used that bijectivity of $T$ implies bijectivity of $T^{\prime\prime}$.
(c) First of all, recall that normality of the cone in an ordered Banach space is equivalent to every order interval being norm-bounded [4, Theorem 2.40]. Now, let $a,b\in Z$. Then $T[a,b]\subseteq[Ta,Tb]$ due to the positivity of $T$ and the later set is norm-bounded because of the normality of the cone in $X$. Hence, $[a,b]=T^{-1}T[a,b]$ is also norm-bounded since $T^{-1}$ is continuous.
For a bipositive operator $J:X\to Z$ between ordered Banach spaces, we are interested in conditions that ensure that $J(X_{+})$ is a face of $Z_{+}$. For this purpose, we introduce the following result regarding extension of linear functionals:
Proposition 2.6.
Let $X$ be an ordered Banach space with a generating cone and let $V$ be a vector subspace of $X$. Let $x^{\prime}\in X^{\prime}_{+}$ and $\varphi:V\to\mathbb{R}$ be linear. The following are equivalent.
[…]
    There exists $y^{\prime}\in X^{\prime}$ such that $0\leq y^{\prime}\leq x^{\prime}$ and $y^{\prime}\raisebox{-2.15277pt}{$|$}_{V}=\varphi$,
[…]
    For all $v\in V$ and $w\in X_{+}$ the inequality $v\leq w$ implies $\varphi(v)\leq\left\langle x^{\prime}\,,\,w\right\rangle$.
Proof 2.7.
“(i) $\Rightarrow$ (ii)”: For each $v\in V,w\in X_{+}$ with $v\leq w$, we have
$\varphi(v)=\left\langle y^{\prime}\,,\,v\right\rangle\leq\left\langle y^{\prime}\,,\,w\right\rangle\leq\left\langle x^{\prime}\,,\,w\right\rangle,$
due to the positivity of $y^{\prime}$. This proves (ii).
“(ii) $\Rightarrow$ (i)”: Let (ii) be true. First of all, note that the functional $p:X\to[0,\infty)$ given by $x\mapsto\inf\{\left\langle x^{\prime}\,,\,w\right\rangle:w\in X_{+},w\geq x\}$ is well-defined since $X_{+}$ is generating and can easily checked to be sublinear. Moreover, one has $\varphi(v)\leq p(v)$ for all $v\in V$ since, if $v\in V$ is given, then $\varphi(v)\leq\left\langle x^{\prime}\,,\,w\right\rangle$ for each $w\in X_{+}$ with $w\geq x$, by assumption. We may thus employ the Hahn-Banach theorem to obtain a linear functional $y^{\prime}:X\to\mathbb{R}$ that satisfies $y^{\prime}\raisebox{-2.15277pt}{$|$}_{V}=\varphi$ and $y^{\prime}(x)\leq p(x)$ for all $x\in X$. For each $x\in X_{+}$, observe that $y^{\prime}(-x)\leq p(-x)=0,$ so $y^{\prime}$ is positive. In particular, $y^{\prime}\in X^{\prime}$ since the cone $X_{+}$ is generating [1, Corollary 6]. Finally,
$\left\langle y^{\prime}\,,\,x\right\rangle\leq p(x)\leq\left\langle x^{\prime}\,,\,x\right\rangle,\quad(x\in X_{+})$
proves that $y^{\prime}\leq x^{\prime}$.
Theorem 2.8.
Let $X$ be an ordered Banach space with a generating and normal cone such that $X_{+}$ is a face of $X^{\prime\prime}_{+}$. Also, let $Z$ be an ordered Banach space such that there exists a bipositive embedding $J:X\to Z$.
Assume that $(R_{n})$ is a sequence in $\mathcal{L}(Z,X)_{+}$ such that $(JR_{n})$ and $(R_{n}J)$ both converge to the identity operator (on $Z$ and $X$ respectively) in the weak operator topology. Then $J(X_{+})$ is a face of $Z_{+}$.
Proof 2.9.
Denoting $k:X\to X^{\prime\prime}$ and $\widetilde{k}:Z\to Z^{\prime\prime}$ by the canonical embeddings, the following diagram commutes:
[…]
Moreover, density of $J$ implies the injectivity of $J^{\prime}:Z^{\prime}\to X^{\prime}$. Let $x\in X_{+}$ and $z\in Z_{+}$ such that $0\leq z\leq Jx$. Define $\varphi:J^{\prime}(Z^{\prime})\to\mathbb{R}$ as
$v\mapsto\varphi(v):=\left\langle\widetilde{k}(z)\,,\,(J^{\prime})^{-1}v\right\rangle.$
Let $v=J^{\prime}z^{\prime}\in J^{\prime}(Z^{\prime})$ and fix $w\in X^{\prime}_{+}$ such that $w\geq v$. Then
$\varphi(v)=\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\left\langle z^{\prime}\,,\,z\right\rangle=\lim_{n\to\infty}\left\langle z^{\prime}\,,\,JR_{n}z\right\rangle=\lim_{n\to\infty}\left\langle v\,,\,R_{n}z\right\rangle.$
Positivity of $z$ and $w$ thus implies that
$\varphi(v)\leq\limsup_{n\to\infty}\left\langle w\,,\,R_{n}z\right\rangle\leq\limsup_{n\to\infty}\left\langle w\,,\,R_{n}Jx\right\rangle=\left\langle w\,,\,x\right\rangle=\left\langle k(x)\,,\,w\right\rangle.$
Thus by Proposition 2.6, there exists $y^{\prime\prime}\in X^{\prime\prime}$ such that $0\leq y^{\prime\prime}\leq k(x)$ and $y^{\prime}\raisebox{-2.15277pt}{$|$}_{J^{\prime}(Z^{\prime})}=\varphi$. Furthermore, for each $z^{\prime}\in Z^{\prime}$, we obtain that
$\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\varphi(J^{\prime}z^{\prime})=\left\langle y^{\prime\prime}\,,\,J^{\prime}z^{\prime}\right\rangle=\left\langle J^{\prime\prime}y^{\prime\prime}\,,\,z^{\prime}\right\rangle$
and consequently, $\widetilde{k}(z)=J^{\prime\prime}y^{\prime\prime}$.
Finally, since $k(X_{+})$ is a face of $X^{\prime\prime}_{+}$ and $0\leq y^{\prime\prime}\leq k(x)$, there exists $y\in X$ such that $y^{\prime\prime}=k(y)$. As $k$ is bipositive, it follows that $0\leq y\leq x$. Wherefore for each $z^{\prime}\in Z^{\prime}$, we have
$\left\langle z^{\prime}\,,\,z\right\rangle=\left\langle\widetilde{k}(z)\,,\,z^{\prime}\right\rangle=\left\langle J^{\prime\prime}k(y)\,,\,z^{\prime}\right\rangle=\left\langle\widetilde{k}(Jy)\,,\,z^{\prime}\right\rangle=\left\langle z^{\prime}\,,\,Jy\right\rangle$
which proves that $z=Jy\in J(X)$, as desired.
We show in Example 2.17 that the assumption $X_{+}$ is a face of $X^{\prime\prime}_{+}$ cannot be dropped in Theorem 2.8.
If $X$ and $D$ are Banach spaces, it is a well-known consequence of the Hahn-Banach extension theorem that every operator $T\in\mathcal{L}(X,D)$ satisfies the norm equality $\left\lVert T^{\prime}\right\rVert_{D^{\prime}\to X^{\prime}}=\left\lVert T\right\rVert_{X\to D}$. The following theorem can be interpreted as a variation of the inequality $\leq$” in this equality. For an intuition about the theorem, we refer to the subsequent Examples 2.1.
Theorem 2.10.
Let $X,D,E$, and $\tilde{E}$ be ordered Banach spaces such that the cone $\tilde{E}_{+}$ is generating and normal. Consider positive bounded linear operators
[…]
such that $\tilde{J}$ is bipositive and its range $\operatorname{Rg}\tilde{J}$ is majorizing in $\tilde{E}$. Then
$\left\lVert(JT)^{\prime}\right\rVert_{\operatorname{span}{E^{\prime}_{+}}\to X^{\prime}}\leq c\;\lVert\tilde{J}T\rVert_{X\to\tilde{E}}$
for a number $c\geq 0$ that might depend on all involved spaces and operators except for $X$ and $T$. Here, $\operatorname{span}(E^{\prime}_{+})=E^{\prime}_{+}-E^{\prime}_{+}$ is endowed with the norm given in formula (1.2) (with $X_{+}$ replaced with $E^{\prime}_{+}$).
Note that the space $D$ does not occur in the conclusion of the theorem. We are interested in the mappings $\tilde{J}T:X\to\tilde{E}$ and $(JT)^{\prime}:E^{\prime}\to X^{\prime}$ and $D$ is an auxiliary space to relate the properties of those two mappings. Before giving the proof of the theorem, we illustrate two situations where the assumptions of the theorem are fulfilled:
{examples}
Let $p\in[1,\infty]$, let $T:L^{p}(\mathbb{R})\to L^{\infty}(\mathbb{R})$ be a positive bounded linear operator, and assume that every function in the range of $T$ is Lipschitz continuous. Let $\operatorname{Lip}(\mathbb{R})$ denote the space of all scalar-valued Lipschitz continuous functions on $\mathbb{R}$, which is an ordered Banach space when endowed with the pointwise order and the norm $\left\lVert f\right\rVert_{\operatorname{Lip}}:=\left\lvert f(0)\right\rvert+\operatorname{Lip}(f)$, where $\operatorname{Lip}(f)$ is the Lipschitz constant of $f$. Then Theorem 2.10 implies that
$\left\lVert T^{\prime}\right\rVert_{\operatorname{span}(\operatorname{Lip}(\mathbb{R})^{\prime}_{+})\to(L^{p})^{\prime}}\leq c\left\lVert T\right\rVert_{L^{p}\to L^{\infty}}$
for a constant $c$ that does not depend on $T$.
To see this, choose $X=L^{p}(\mathbb{R})$, $E=\operatorname{Lip}(\mathbb{R})$, $\tilde{E}=L^{\infty}(\mathbb{R})$ and $D=E\cap\tilde{E}$ with the norm $\left\lVert f\right\rVert_{D}:=\left\lVert f\right\rVert_{\operatorname{Lip}}+\left\lVert f\right\rVert_{\infty}$ in the theorem, let $J$ and $\tilde{J}$ be the canonical embeddings, and observe that $T$ is bounded from $L^{p}(\mathbb{R})$ to $D$ by the closed graph theorem.
(b) Let $\Omega$ denote the open unit ball in $\mathbb{R}^{d}$ and the consider Banach lattices $X=L^{2}(\Omega)$ and $\widetilde{E}=L^{\infty}(\Omega)$. Let $E$ and $D$ be ordered Banach spaces $H^{4}(\Omega)\cap H_{0}^{2}(\Omega)$ and $E\cap\widetilde{E}$ respectively, where the latter is equipped with the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{D}:=\left\lVert\mathord{\,\cdot\,}\right\rVert_{E}+\left\lVert\mathord{\,\cdot\,}\right\rVert_{\infty}$; note that the inclusion $E\subseteq\tilde{E}$ holds if $d<8$, but not otherwise.
As $D$ contains the constant one function $\operatorname{{\mathbbm{1}}}$, so $D$ is majorizing in $\widetilde{E}$. In particular, the canonical embeddings of $D$ into $E$ and $\widetilde{E}$ satisfy the assumptions of Theorem 2.10. Moreover, the bi-Laplace operator $A:=-\Delta^{2}$ with domain $E$ is the generator of an analytic uniformly eventually positive analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $X$ by [13, Theorem 4.4], i.e., there exists $t_{0}\geq 0$ such that $T(t)\geq 0$ for all $t\geq t_{0}$. Furthermore, $\operatorname{dom}A^{n}\subseteq\widetilde{E}$ for sufficiently large $n\in\mathbb{N}$, so $T(t)X\subseteq D$ for all $t>0$. Theorem 2.10 thus ensures that there exists $c>0$ such that
$\left\lVert T(t)^{\prime}\right\rVert_{\operatorname{span}{E^{\prime}_{+}}\to L^{2}(\Omega)}\leq c\left\lVert T(t)\right\rVert_{L^{2}(\Omega)\to L^{\infty}(\Omega)}$
for all $t\geq t_{0}$.
Proof 2.11 (Proof of Theorem 2.10).
As $\tilde{E}_{+}$ is generating cone and $\tilde{J}(D)$ is majorizing in $\tilde{E}$, one can apply Proposition 2.1(c) to get $\tilde{J}^{\prime}(\tilde{E}^{\prime}_{+})=D^{\prime}_{+}$. On the other hand, the dual cone $\tilde{E}^{\prime}_{+}$ is also generating owing to the normality of the cone $\tilde{E}_{+}$ [4, Theorem 2.26] and thus $\tilde{J}^{\prime}\tilde{E}^{\prime}=D^{\prime}_{+}-D^{\prime}_{+}=\operatorname{span}D^{\prime}_{+}$. We now endow $\operatorname{span}D^{\prime}_{+}$ with the norm given in formula (1.2) (with $X_{+}$ replaced with $D^{\prime}_{+}$) and note that $\tilde{J}^{\prime}$ is continuous from $\tilde{E}^{\prime}$ to $\operatorname{span}D^{\prime}_{+}$ due to the closed graph theorem. Hence, the open mapping theorem implies that there exists a set $S\subseteq\tilde{E}^{\prime}$ that is norm-bounded by a number $\tilde{c}\geq 0$ such that $\tilde{J}^{\prime}S$ covers the unit ball in $\operatorname{span}D^{\prime}_{+}$. As a result,
$\left\lVert T^{\prime}\right\rVert_{\operatorname{span}D^{\prime}_{+}\to X^{\prime}}\leq\tilde{c}\left\lVert T^{\prime}\tilde{J}^{\prime}\right\rVert_{\tilde{E}^{\prime}\to X^{\prime}}=\tilde{c}\left\lVert\tilde{J}T\right\rVert_{X\to\tilde{E}}.$
On the other hand, as $J^{\prime}$ is positive it maps $\operatorname{span}E^{\prime}_{+}$ into $\operatorname{span}D^{\prime}_{+}$ and it is a continuous operator between those Banach spaces by the closed graph theorem. Thus,
$\left\lVert(JT)^{\prime}\right\rVert_{\operatorname{span}\tilde{E}^{\prime}_{+}\to X^{\prime}}\leq\left\lVert T^{\prime}\right\rVert_{\operatorname{span}D^{\prime}_{+}\to X^{\prime}}\left\lVert J^{\prime}\right\rVert_{\operatorname{span}E^{\prime}_{+}\to\operatorname{span}D^{\prime}_{+}},$
which combined with the previous estimate completes the proof.
2.2 The order on $X_{-1}$
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on an ordered Banach space $X$. Throughout we endow the extrapolation space $X_{-1}$ with the cone
$X_{-1,+}:=\overline{X_{+}}^{\left\lVert\mathord{\,\cdot\,}\right\rVert_{X_{-1}}},$
i.e., the norm closure of $X_{+}$ in $X_{-1}$, and consider the partial order that is induced by $X_{-1,+}$. This order on $X_{-1}$ was, for the case where $X$ is a Banach lattice, introduced in [10] in the context of perturbation theorems for positive semigroups. A number of fundamental properties of $X_{-1,+}$ – in particular that it is indeed a cone – are proved in [10, Remark 2.2 and Proposition 2.3]. We now show that the same properties remain true on ordered Banach spaces. Most arguments are similar, but we include the details for the convenience of the reader. Our proof of the equality $X_{-1,+}\cap(-X_{-1,+})=\{0\}$ is a bit different since the cone $X_{+}$ need not be normal in the following proposition.
Proposition 2.12.
Let $X$ be an ordered Banach space and let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on $X$.
[…]
    The set $X_{-1,+}$ is a cone and hence, $(X_{-1},X_{-1,+})$ is an ordered Banach space.
[…]
    One has $X_{+}=X_{-1,+}\cap X$. In other words, two vectors $x,y\in X$ satisfy $x\leq y$ in $X$ if and only if they satisfy the same inequality in $X_{-1}$.
[…]
    For every $\lambda>s(A)$, the resolvent operator $R(\lambda,A_{-1})$ is positive from $X_{-1}$ to $X$.
Proof 2.13.
(c) Let $\lambda>s(A)=s(A_{-1})$. We prove that $R(\lambda,A_{-1})X_{-1,+}\subseteq X_{+}$ (this is precisely the claim of (c), but we are only allowed to speak of a positive operator from $X_{-1}$ to $X$ once we have proved (a)). So let $x\in X_{-1,+}$ and choose a sequence $(x_{n})\subseteq X_{+}$ that converges to $x$ in $X_{-1}$. Firstly, if $\lambda>\omega_{0}(A)$, then the operator $R(\lambda,A)$ is positive from $X$ to $X$ owing to the Laplace transform representation of the resolvent. Now, the positivity extends from the interval $(\omega_{0}(A),\infty)$ to the interval $(s(A),\infty)$ via the Taylor expansion of the resolvent and a connectedness argument; see for instance [23, Proposition 2.1(a)] for details. Therefore, $R(\lambda,A)x_{n}\in X_{+}$ for each index $n$. As $R(\lambda,A_{-1})$ is continuous from $X_{-1}$ to $X$, it follows that
$R(\lambda,A_{-1})x=\lim_{n\to\infty}R(\lambda,A_{-1})x_{n}=\lim_{n\to\infty}R(\lambda,A)x_{n}\in X_{+},$
where both limits are taken in the Banach space $X$.
(a) It is easy to see that $X_{-1,+}$ is closed, convex, and that $\alpha X_{-1,+}\subseteq X_{-1,+}$ for each $\alpha\in[0,\infty)$, so one only has to show that $X_{-1,+}\cap-X_{-1,+}=\{0\}$. To this end, suppose $x\in X_{-1}$ satisfies ±$\pm x\geq 0$. Fixing $\lambda>s(A)$, it follows from (c) that ±$\pm R(\lambda,A_{-1})x\in X_{+}$. Since $X_{+}$ is a cone, this implies that $R(\lambda,A_{-1})x=0$, and so $x=0$ by injectivity of the resolvent operator.
(b) Obviously, $X_{+}\subseteq X_{-1,+}\cap X$. To establish the converse inclusion, we let $x\in X_{-1,+}\cap X$. It follows from (c) that $\lambda R(\lambda,A)x=\lambda R(\lambda,A_{-1})x$ is in $X_{+}$ for all $\lambda>\max\{s(A),0\}$. Moreover, since $x$ is in $X$ one has $\lambda R(\lambda,A)x\to x$ in $X$ as $\lambda\to\infty$. As $X_{+}$ is closed in $X$, we even get that $x\in X_{+}$.
It is clear that the extrapolated semigroup $(T_{-1}(t))_{t\geq 0}$ on $X_{-1}$ leaves $X_{-1,+}$ invariant. Since $X_{-1}$ is, by the previous proposition, an ordered Banach space with respect to this cone, we can rephrase this by saying that the extrapolated semigroup is also positive.
Even if $X$ is a Banach lattice, the cone $X_{-1,+}$ is, in general, not generating in $X_{-1}$ [10, Examples 5.1 and 5.3]. In particular, $X_{-1}$ is usually not a Banach lattice. Nevertheless, for any $\lambda>s(A)$, we know from Proposition 2.12 that the operator $R(\lambda,A_{-1}):X_{-1}\to\operatorname{dom}A_{-1}=X$ is a continuous and positive linear bijection. This allows us to apply the results proved in Section 2.1 to $Z=X_{-1}$ in order to obtain various order properties of $X_{-1}$. First of all, Theorem 2.4 immediately implies the following (cf. Table 1):
Corollary 2.14.
Let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on an ordered Banach space $X$. The following are true.
[…]
    If every positive, increasing, and norm-bounded net in $X$ is norm-convergent, then the same is true in $X_{-1}$.
[…]
    If the cone of $X$ is the face of the bidual cone, then the same is true for the cone of $X_{-1}$.
[…]
    If the cone of $X$ is normal, then so is the cone of $X_{-1}$.
As a consequence of Theorem 2.8, we get sufficient conditions for the cone $X_{+}$ to be a face of the cone $X_{-1,+}$.
Corollary 2.15.
Let $X$ be an ordered Banach space with a generating and normal cone and assume that $X_{+}$ is a face of $X^{\prime\prime}_{+}$. If $(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$, then $X_{+}$ is a face of $X_{-1,+}$.
Proof 2.16.
Fix $\lambda>s(A)$. As pointed out before, $T=R(\lambda,A_{-1})$ is a continuous and positive linear bijection from $X_{-1}$ to $X$. Moreover, by definition $X$ is dense in $X_{-1}$ and the equality $X_{+}=X_{-1,+}\cap X$ from Proposition 2.12 implies the canonical embedding is bipositive. Therefore, defining $R_{n}:=nR(n,A_{-1})$ for $n\in\mathbb{N}$, Theorem 2.8 implies that $X_{+}$ is indeed a face of $X_{-1,+}$.
The following example shows that in general $X_{+}$ need not be a face in $X_{-1,+}$.
Example 2.17.
On the space $E=\{f\in C[0,1]:f(0)=f(1)\}$, the operator
$\operatorname{dom}A:=\{f\in C^{1}[0,1]\in E:f^{\prime}(0)=f^{\prime}(1)\},\qquad f\mapsto f^{\prime}$
generates a positive periodic shift semigroup [41, Section A-I.2.5] for which the extrapolation space was shown in [10, Example 5.3] to be
$X_{-1}=\{g\in\mathcal{D}(0,1)^{\prime}:g=f-\partial f\text{ for some }f\in X\};$
here $\mathcal{D}(0,1)$ is the space of test functions on $(0,1)$ and $\partial$ denotes the distributional derivative. The function
$f(x)=\begin{cases}1-\frac{e^{x}}{1+\sqrt{e}}\qquad&\text{ for }x\in[0,\frac{1}{2})\\ \frac{e^{x}}{e+\sqrt{e}}\qquad&\text{ for }x\in[\frac{1}{2},1]\end{cases}$
lies in $X$ and satisfies $f-\partial f=\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$. In particular, the characteristic function $\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$ lies in $X_{-1,+}$ and satisfies $0\leq\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}\leq\operatorname{{\mathbbm{1}}}$. As $X_{+}$ contains $\operatorname{{\mathbbm{1}}}$ but not $\operatorname{{\mathbbm{1}}}_{[0,\frac{1}{2}]}$, it follows that $X_{+}$ is not a face of $X_{-1,+}$.
We now present an application of Theorem 2.10 and follow it up with two examples involving elliptic operators. To state the result, let us first recall the definition of Favard spaces with index $\beta\in(-1,0]$. Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$, fix an arbitrary number $\omega$ that is strictly larger than the growth bound of the semigroup, and let $\beta\in(-1,0]$. For each $x\in X_{-1}$ one defines
$\left\lVert x\right\rVert_{\beta}:=\sup_{t\in(0,\infty)}\frac{1}{t^{\beta+1}}\left\lVert e^{-\omega t}T_{-1}(t)x-x\right\rVert_{-1}\in[0,\infty].$
The space $F_{\beta}:=\{x\in X_{-1}:\,\left\lVert x\right\rVert_{\beta}<\infty\}$ is called the Favard space of order $\beta$ of the semigroup. It is a Banach space with respect to the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{\beta}$; see, for instance, [17, Definition II.5.10 and II.5.11] for details.
Theorem 2.18.
Let $X$ be a reflexive ordered Banach space, let $\varphi\in X^{\prime}$ be strictly positive, and let $(T(t))_{t\geq 0}$ be a positive and immediately differentiable $C_{0}$-semigroup. Assume that $T(t)^{\prime}X^{\prime}\subseteq(X^{\prime})_{\varphi}$ for each $t>0$ (where the definition of $(X^{\prime})_{\varphi}$ is given in (1.3)) and that $\operatorname{dom}A^{\prime}\cap(X^{\prime})_{\varphi}$ is majorizing in $(X^{\prime})_{\varphi}$.
[…]
    There exists a number $c\geq 0$ such that
    $\left\lVert T_{-1}(t)\right\rVert_{\operatorname{span}X_{-1,+}\to X}\leq c\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}$
    for all $t>0$.
[…]
    If $(T(t))_{t\geq 0}$ is additionally analytic and the dual semigroup satisfies the ultracontractivity type estimate $\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\leq\widetilde{c}t^{-\alpha}$ for constants $\widetilde{c}\geq 0$ and $\alpha\in[0,1)$ and for all $t\in(0,1]$, then $\operatorname{span}X_{-1,+}$ is contained in the negative Favard space $F_{-\alpha}$ associated to $(T(t))_{t\geq 0}$.
Proof 2.19.
(a) We apply Theorem 2.10 to the following situation: choose the space $X$ from Theorem 2.10 as the space $X^{\prime}$ in the present theorem, let $E:=\operatorname{dom}A^{\prime}$, $\widetilde{E}:=(X^{\prime})_{\varphi}$, and $D:=E\cap\widetilde{E}$ which is an ordered Banach space with respect to the norm $\left\lVert\mathord{\,\cdot\,}\right\rVert_{D}:=\left\lVert\mathord{\,\cdot\,}\right\rVert_{E}+\left\lVert\mathord{\,\cdot\,}\right\rVert_{\widetilde{E}}$. Let $J$ and $\widetilde{J}$ be the canonical embeddings of $D$ into $E$ and $\widetilde{E}$ and note that the range of $\widetilde{J}$ is majorizing in $\widetilde{E}$ by the assumptions of the present theorem. Fix $t>0$ and let $T:=T(t)^{\prime}$ – our assumptions imply that $\operatorname{Rg}T\subseteq D$. Hence, Theorem 2.10 yields a $c>0$, independent of $t$ such that
$\left\lVert(T(t)^{\prime})^{\prime}\right\rVert_{\operatorname{span}({\operatorname{dom}(A^{\prime})^{\prime}_{+}})\to X^{\prime\prime}}\leq c\;\lVert\widetilde{T}(t)^{\prime}\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}.$
Due to the reflexivity, one can replace the left-hand side of the estimate with the norm $\left\lVert T_{-1}(t)\right\rVert_{\operatorname{span}X_{-1,+}\to X}$, whence the claim.
(b) This is a consequence of (a) and a characterisation [17, Proposition II.5.13 and Definition II.5.11] of Favard spaces for analytic semigroups.
Example 2.20 ($X_{-1}$ for elliptic operators with Neumann boundary conditions).
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary and let $p\in(1,\infty)$. Let $a\in L^{\infty}(\Omega;\mathbb{R}^{d\times d})$ satisfy the following coercivity condition: there exists a number $\nu>0$ such that for almost all $x\in\Omega$ the estimate
$\xi^{T}a(x)\overline{\xi}\geq\nu\left\lVert\xi\right\rVert^{2}\qquad\text{for all }\xi\in\mathbb{C}^{d}$
holds. Consider the divergence form elliptic operator $A:L^{p}(\Omega)\supseteq\operatorname{dom}A\to L^{p}(\Omega)$, $u\mapsto\operatorname{div}(a\nabla u)$ with Neumann boundary conditions. This operator can be constructed using form methods in $L^{2}(\Omega)$ and then extrapolating to the $L^{p}$-scale, see [44, Section 4.1]. Let $q\in(1,\infty)$ be the Hölder conjugate of $p$. The operator $A$ generates a positive analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $L^{p}(\Omega)$ [44, Corollary 4.3 and Theorem 1.52]. Further, if $\alpha:=\frac{d}{2q}<1$, then the following hold.
  1. (a)
    The span of the cone $L^{p}(\Omega)_{-1,+}$ is contained in the negative Favard space $F_{-\alpha}$.
  2. (b)
    The cone $L^{p}(\Omega)_{-1,+}$ consists precisely of all positive finite Borel measures on $\overline{\Omega}$.
Let us show how the claims (a) and (b) follow from the theory developed so far.
(a) The dual operator $A^{\prime}$ on the Banach space $L^{q}(\Omega)$ is also a divergence form elliptic operator with Neumann boundary conditions, but with the diffusion coefficient $a^{T}$. Therefore the semigroup $(T^{\prime}(t))_{t\geq 0}$ generated by $A^{\prime}$ maps $L^{q}$ into $L^{\infty}(\Omega)=\big{(}L^{q}(\Omega)\big{)}_{\operatorname{{\mathbbm{1}}}}$ and satisfies the ultracontractivity estimate
$\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to L^{\infty}}\leq ct^{-\frac{d}{2}\frac{1}{q}}=ct^{-\alpha}$
for some $c>0$ and all $t\in(0,1]$; see [6, Sections 7.3.2 and 7.3.6]. Moreover, $\operatorname{dom}A^{\prime}$ contains the constant function $\operatorname{{\mathbbm{1}}}$, so $\operatorname{dom}A^{\prime}\cap L^{\infty}(\Omega)$ is majorizing in $L^{\infty}(\Omega)$. Hence, the assumptions of Theorem 2.18(b) are satisfied.
(b) As $L^{p}(\Omega)$ is reflexive, the space $L^{p}(\Omega)_{-1}$ can be identified with $\operatorname{dom}(A^{\prime})^{\prime}$ as an ordered Banach space. So we only need to identify the positive linear functionals on $\operatorname{dom}A^{\prime}$. Since $\alpha<1$ it follows that $\operatorname{dom}A^{\prime}$ is contained in the space $C(\overline{\Omega})$ [43, Proposition 3.6]. Moreover, $\operatorname{dom}A^{\prime}$ is majorizing in $C(\overline{\Omega})$ since it contains the constant functions and it is dense in $C(\overline{\Omega})$ according to [43, Lemma 4.2]. Since the cone in $C(\overline{\Omega})$ has non-empty interior it follows that even the cone $\operatorname{dom}A^{\prime}_{+}$ is dense in the cone $C(\overline{\Omega})$. So by Proposition 2.1(b) the dual space $C(\overline{\Omega})^{\prime}$ embeds bipositively into the space $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}(\Omega)_{-1}$ and by part (c) of the same proposition, this embedding maps the cone of $C(\overline{\Omega})^{\prime}$ surjectively onto the cone of $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}(\Omega)_{-1}$. Since the positive cone in $C(\overline{\Omega})^{\prime}$ consists of all positive finite Borel measures on $\overline{\Omega}$, the claim follows.
For elliptic operators with Dirichlet boundary conditions, the situation is a bit more subtle: in order to apply Theorem 2.18 one needs the semigroup to be intrinsically ultracontractive, which means that it not only maps $L^{1}$ into $L^{\infty}$ but also maps a weighted $L^{1}$-space – with the leading eigenfuntion $u$ of the operator as a weight – into the principal ideal generated by $u$. Such results are available in the literature in the case that the domain and the coefficients of the differential operator are sufficiently smooth, see for instance, [14, Theorem 4.6.2]. To get an estimate from $L^{p}$ into the principal ideal generated by $u$ we need $p$ to be at least $2$ – this is why the result in the following example can only be applied if the dimension of the spatial domain is low enough.
Example 2.21 ($X_{-1}$ for elliptic operators with Dirichlet boundary conditions).
Let $p\in(1,\infty)$ and let $A:L^{p}(\Omega)\supseteq\operatorname{dom}A\to L^{p}(\Omega)$ denote the same operator as in Example 2.20, but now with Dirichlet boundary conditions. Assume in addition that the bounded domain $\Omega\subseteq\mathbb{R}^{d}$ has $C^{2}$-boundary and that the coefficient matrix $a$ is symmetric and $C^{1}$ on $\overline{\Omega}$. Again, it is well-known that $A$ generates a positive and analytic $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on $L^{p}(\Omega)$.
Assume now that $p\geq 2$ and that $\alpha:=\frac{1}{2q}(1+\frac{d}{2})<1$ (note that both conditions can only be simultaneously satisfied if $d\leq 5$); here $q$ denotes the Hölder conjugate of $p$. Then $L^{p}(\Omega)_{-1,+}$ is contained in the negative Favard space $F_{-\alpha}$ and thus in a real interpolation space between $X$ and $X_{-1}$.
To prove this, first note that since the coefficient matrix $a$ is assumed to be symmetric, the dual operator $T(t)^{\prime}$ acts as the continuous extension of $T(t)$ to the space $L^{q}(\Omega)$ for the Hölder conjugate $q$ of $p$. Let $u\in L^{p}(\Omega)_{+}$ denote the leading eigenfunction of $A$ and thus also of $A^{\prime}$ (so one also has $u\in L^{q}(\Omega)_{+}$) associated to the leading eigenvalue $-\lambda_{0}\in(-\infty,0)$. We show that the assumptions of Theorem 2.18(a) are satisfied for the functional $\varphi:=u\in L^{q}(\Omega)$.
The smoothness assumptions on $\partial\Omega$ and on the coefficient matrix $a$ imply that the heat kernel $k_{t}:\Omega\times\Omega\to[0,\infty)$ of the semigroup operator $T(t)$ (and thus also of $T(t)^{\prime}$) satisfies the intrinsic ultracontractivity estimate
(2.1) $k_{t}(x,y)\leq ct^{-\frac{1}{2}(1+\frac{d}{2})}u(x)u(y)=ct^{-\alpha q}u(x)u(y)$
for a constant $c>0$, all $t>0$, and all $x,y\in\Omega$ [14, Theorem 4.6.2] (it should be noted that there is a typo in the formula in [14, Theorem 4.6.2]: the factor $-1$ in the exponent should actually be $-\frac{1}{2}$, as the proof shows). Hence, $T(t)^{\prime}$ maps each $L^{q}(\Omega)$ into the principal ideal $(L^{q}(\Omega))_{u}$. Moreover, since $u\in\operatorname{dom}A^{\prime}$, the space $\operatorname{dom}A^{\prime}\cap(L^{q}(\Omega))_{u}$ is clearly majorizing in $(L^{q}(\Omega))_{u}$. So it only remains to show that $\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to(L^{q})_{u}}\leq\tilde{c}t^{-\alpha}$ for a constant $\tilde{c}>0$ and all $t\in(0,1]$.
To this end, we first observe that the semigroup $(T(t))_{t\geq 0}$ is dominated by the semigroup generated by the same differential operator but with Neumann boundary conditions; this follows for instance from [44, Corollary 2.22]. In addition to the principal ideals $(L^{p}(\Omega))_{u}$ and $(L^{q}(\Omega))_{u}$ we will now use the weighted $L^{1}$-space $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$. This space coincides with the norm completion of $L^{p}(\Omega)$ with respect to the norm $\left\lVert f\right\rVert_{L^{1}(\mathrm{d}x)}:=\int_{\Omega}\left\lvert f\right\rvert u\;\mathrm{d}x$ and behaves, in a sense, dually to the principal ideal $(L^{q}(\Omega))_{u}$; we refer to [9, Section 2] or [13, Section 2] for a detailed discussion.
Moreover, since $T(t)^{\prime}u=e^{-t\lambda_{0}}u$ for each $t$, the operator $T(t)$ extends to a bounded linear operator on $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$, which we again denote by $T(t)$ and which satisfies
$\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)}\leq e^{-t\lambda_{0}}\leq 1.$
At the same time it follows from the heat kernel estimate (2.1) that, for each $t>0$, the operator $T(t)$ maps $L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$ into the principal ideal $(L^{p}(\Omega))_{u}$ with norm
$\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to(L^{p})_{u}}\leq ct^{-\alpha q}$
(an abstract version of this argument can be found in [9, Proposition 2.2]). We will now combine the previous two norm estimates for $T(t)$ with an interpolation inequality to get an estimate for $\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to L^{p}}$: for every function $f\in(L^{p}(\Omega))_{u}\subseteq L^{p}(\Omega)\subseteq L^{1}(\Omega,u\nobreak\hskip 1.00006pt\mathrm{d}x)$ we have
$\left\lVert f\right\rVert_{L^{p}}^{p}=\int_{\Omega}u^{p-2}\;\left\lvert f\right\rvert u\;\frac{\left\lvert f\right\rvert^{p-1}}{u^{p-1}}\;\mathrm{d}x\leq\left\lVert u^{p-2}\right\rVert_{\infty}\left\lVert f\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)}\left\lVert f\right\rVert_{(L^{p})_{u}}^{p-1},$
so $\left\lVert f\right\rVert_{L^{p}}\leq\left\lVert u^{p-2}\right\rVert_{\infty}^{1/p}\left\lVert f\right\rVert_{L^{1}(u\;\mathrm{d}x)}^{1/p}\left\lVert f\right\rVert_{(L^{p})_{u}}^{1/q}$; here we have also used that $u\in L^{\infty}(\Omega)$ and the assumption $p\geq 2$, which ensures that $\left\lVert u^{p-2}\right\rVert_{\infty}<\infty$. This estimate for $\left\lVert f\right\rVert_{L^{p}}$ together with the two aforementioned norm estimates for $T(t)$ readily gives the interpolated norm estimate
$\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to(L^{p})_{u}}\leq c^{1/q}\left\lVert u^{p-2}\right\rVert_{\infty}^{1/p}t^{-\alpha}$
for all $t>0$. For every $t>0$, one has $\left\lVert T(t)\right\rVert_{L^{1}(u\nobreak\hskip 0.70004pt\mathrm{d}x)\to(L^{p})_{u}}=\left\lVert T(t)^{\prime}\right\rVert_{L^{q}\to(L^{q})_{u}}$. Indeed, the inequality $\leq$ is shown in an abstract setting in [13, Proposition 2.1], but one can show that even equality is true by using the duality result in [47, Exercise IV.9(a)]. So the assumption of Theorem 2.18(b) is satisfied, which gives the claimed result.
For Dirichlet boundary conditions, not all elements of $L^{p}(\Omega)_{-1,+}$ can be described as measures on $\overline{\Omega}$. This is in contrast to the situation for Neumann boundary conditions and small $p$, see Example 2.20(b). One can see this easily by taking $A$ to be the Dirichlet Laplacian on the interval $(-1,1)$. Then $A^{\prime}$ is also the Dirichlet Laplacian and the domain $\operatorname{dom}A^{\prime}$ is (no matter which value we use for $p$) a majorizing and dense subset of $C^{1}[-1,1]$. Hence, the linear functionals $f\mapsto f^{\prime}(-1)$ and $f\mapsto-f^{\prime}(1)$ are positive elements of $\operatorname{dom}(A^{\prime})^{\prime}\simeq L^{p}((-1,1))_{-1}$ that are not given by measures on $[-1,1]$.
3 Admissibility of observation operators
We now come to the main topic of the paper – admissibility. Throughout this section, we consider the linear time-invariant system
$\Sigma(A,C)\quad\left\{\begin{aligned} \dot{x}(t)&=Ax(t),\quad&t\geq 0\\ y(t)&=Cx(t),\quad&t\geq 0\\ x(0)&=x_{0}\end{aligned}\right.;$
where $A$ is the generator of a $C_{0}$-semigroup $(T(t))_{t\geq 0}$ on a Banach space $X$ and the observation operator $C\in\mathcal{L}(X_{1},Y)$ takes values in a Banach space $Y$. The output map is defined as
(3.1) $\Psi_{\tau}:X_{1}\to L^{1}([0,\tau],Y),\qquad x\mapsto CT(\mathord{\,\cdot\,})x$
for a fixed $\tau>0$. We say that $C$ is a $\mathrm{L}^{1}$-admissible observation operator if $\Psi_{\tau}$ has a bounded extension to $X$ for some (equivalently, all) $\tau>0$. This is of course, equivalent to the existence of $K_{\tau}>0$ such that
(3.2) $\left\lVert CT(\mathord{\,\cdot\,})x\right\rVert_{L^{1}([0,\tau],Y)}\leq K_{\tau}\left\lVert x\right\rVert,\qquad(x\in X_{1}).$
In addition, if we have $\limsup_{\tau\downarrow 0}\left\lVert\Psi_{\tau}\right\rVert_{\mathcal{L}(X,\mathrm{L^{1}}([0,\tau],Y))}=0$, then $C$ is called a zero-class $\mathrm{L^{1}}$-admissible observation operator. Most of our results in this section and the sequel are proved in the following setting:
{assumption}
Suppose that $X$ is an ordered Banach space with a generating and normal cone and $(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$. Moreover, $Y$ is a Banach space and $C\in\mathcal{L}(X_{1},Y)$.
In all of our results, the order on $X_{1}$ is assumed to be the order which it inherits from $X$. We begin with obtaining sufficient conditions for an operator $C\in\mathcal{L}(X_{1},Y)$ to be $L^{1}$-admissible. Thereafter, in Section 3.2, we look at situations where $L^{1}$-admissibility of $C$ automatically implies zero-class $L^{1}$-admissibility.
3.1 $L^{1}$-admissibility
For positive semigroups on Banach lattices, we know from [56, Theorem 4.1.17] that positive observation operators mapping into AL-spaces are always $L^{1}$-admissible. In our first result, we generalise this to ordered Banach spaces whose norm is additive on the positive cone.
Proposition 3.1.
Suppose that Assumption 3 holds. Assume further that $Y$ is an ordered Banach space whose norm is additive on the positive cone and that $C\in\mathcal{L}(X_{1},Y)$ is positive. Then $C$ is $L^{1}$-admissible.
For the proof of Proposition 3.1, we essentially adapt the argument of [56, Theorem 4.1.17] to our setting. The essential ingredient of the argument is that due to the positivity of the semigroup and the observation operator, it suffices to establish (3.2) for positive $x\in X_{1}$. When $X$ and $Y$ are both Banach lattices, this was proved in [19, Lemma 3.1]. However, the result remains true under our assumption as well and is actually a consequence of the following technical lemma.
Lemma 3.2.
Let $X$ be an ordered Banach space with a generating cone, let $\widetilde{X}$ be densely embedded in $X$, and let $Y$ be a Banach space. Let $(F(t))_{t\geq 0}$ be a family of operators in $\mathcal{L}(\widetilde{X},Y)$ with the following properties:
[…]
    If $x\in\widetilde{X}$ and $y,z\in X_{+}$ are such that $x=y-z$, then there exist sequences $(y_{n})$ and $(z_{n})$ in $\widetilde{X}_{+}$ such that $y_{n}\to y$ and $z_{n}\to z$ in $X$ and $\left\lVert F(t)(y_{n}-z_{n})\right\rVert\to\left\lVert F(t)x\right\rVert$ for every $t\geq 0$.
[…]
    The mapping $t\mapsto F(\mathord{\,\cdot\,})x$ lies in $L^{1}([0,\tau],Y)$ for each $x\in\widetilde{X}$.
[…]
    There exists $\widetilde{M}>0$ such that $\int_{0}^{\tau}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq\widetilde{M}\left\lVert x\right\rVert$ for all $x\in\widetilde{X}_{+}$.
Then there exists $M>0$ such that
$\int_{0}^{\tau}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq M\left\lVert x\right\rVert$
for all $x\in\widetilde{X}$.
Proof 3.3.
Let $x\in\widetilde{X}$. By the uniform decomposition property of ordered Banach spaces with generating cones [4, Theorem 2.37], there exists $c>0$ – independent of $x$ – and $y,z\in X_{+}$ such that $x=y-z$ and $\left\lVert y\right\rVert,\left\lVert z\right\rVert\leq c\left\lVert x\right\rVert$. Let $(y_{n})$ and $(z_{n})$ be sequences as in assumption (a). Then
$\int_{0}^{\tau}\left\lVert F(t)(y_{n}-z_{n})\right\rVert\;\mathrm{d}t\leq\widetilde{M}\left(\left\lVert y_{n}\right\rVert+\left\lVert z_{n}\right\rVert\right)$
for each $n\in\mathbb{N}$. An application of the dominated convergence theorem now implies that
$\int_{0}^{\tau}\left\lVert F(t)x\right\rVert\;\mathrm{d}t\leq\widetilde{M}(\left\lVert y\right\rVert+\left\lVert z\right\rVert)\leq 2c\widetilde{M}\left\lVert x\right\rVert,$
as desired.
Proof 3.4 (Proof of Proposition 3.1).
Without loss of generality, assume that $\omega_{0}(A)<0$ and let $C\in\mathcal{L}(X_{1},Y)$ be positive. Then for every positive $x\in X_{1}$, we have
[…]
here the first equality holds because the norm on the positive cone of $Y$ is additive.
On the other hand, if $x\in X_{1}$, there exist $y,z\in X_{+}$ such that $x=y-z$ because the cone of $X$ is generating. Define $y_{n}:=nR(n,A)y$ and $z_{n}:=nR(n,A)z$. Then $(y_{n})$ and $(z_{n})$ are positive sequences in $X_{1}$ converging respectively to $y$ and $z$ in $X$ such that for each $t\geq 0$, the sequence $x_{n}:=y_{n}-z_{n}$ satisfies
$\Big{\lvert}\left\lVert CT(t)x_{n}\right\rVert-\left\lVert CT(t)x\right\rVert\Big{\rvert}$ $=\Big{\lvert}\left\lVert CA^{-1}T(t)Ax_{n}\right\rVert-\left\lVert CA^{-1}T(t)Ax\right\rVert\Big{\rvert}$
$\leq\left\lVert CA^{-1}\right\rVert\left\lVert T(t)\right\rVert\left\lVert Ax_{n}-Ax\right\rVert\to 0$
as $n\to\infty$. The assertion is now an immediate consequence of Lemma 3.2.
The following sufficient criterion for $L^{1}$-admissibility of observation operators follows directly from Proposition 3.1.
Theorem 3.5.
Suppose that Assumption 3 holds. Assume further that there exists an ordered Banach space $\widetilde{X}$ whose norm is additive on the cone $\widetilde{X}_{+}$ such that $C$ factorises as
$C:X_{1}\overset{C_{1}}{\longrightarrow}\widetilde{X}\overset{C_{2}}{\longrightarrow}Y$
for a positive operator $C_{1}\in\mathcal{L}(X_{1},\widetilde{X})$ and an operator $C_{2}\in\mathcal{L}(\widetilde{X},Y)$. Then $C$ is $L^{1}$-admissible.
Proof 3.6.
By Proposition 3.1, the operator $C_{1}$ is $L^{1}$-admissible. Thus there exists $K_{\tau}>0$ such that
$\left\lVert C_{2}C_{1}T(t)x\right\rVert_{L^{1}([0,\tau],Y)}\leq K_{\tau}\left\lVert C_{2}\right\rVert\left\lVert x\right\rVert$
for all $x\in X_{1}$, As a result, $C=C_{2}C_{1}$ is also $L^{1}$-admissible.
The factorisation condition in the preceding corollary might seem a bit artificial at first glance. However, we now give two corollaries which show that the condition is quite natural in some situations.
Corollary 3.7.
Suppose that Assumption 3 holds. If $Y$ is an ordered Banach space and $C\in\mathcal{L}(X_{1},Y)$ is a positive finite-rank operator, then $C$ is $L^{1}$-admissible.
Proof 3.8.
To apply Theorem 3.5, we define $\widetilde{X}\subseteq Y$ to be the span of the range $\operatorname{Rg}C$ and we endow it with the cone $\widetilde{X}_{+}:=\widetilde{X}\cap Y_{+}$. Then $\widetilde{X}$ is a finite-dimensional ordered Banach space and hence, there exists an equivalent norm on this space that is additive on the positive cone; we refer to [4, Corollary 3.8] and [57, Theorem VII.1.1]. So if we set $C_{1}:=C$ (but with co-domain $\widetilde{X}$) and let $C_{2}$ be the canonical embedding of $\widetilde{X}$ into $Y$, Theorem 3.5 is applicable.
One can also use the factorisation assumption in Theorem 3.5 in the following way: the dual statement of this assumption can be characterised in terms of ordered boundedness, as we show in Appendix A. If $Y$ is reflexive, one can thus reformulate Theorem 3.5 in the following way.
Corollary 3.9.
Suppose that Assumption 3 holds and assume that $Y$ is reflexive. If the dual operator $C^{\prime}:Y^{\prime}\to(X_{1})^{\prime}=(X^{\prime})_{-1}$ maps the unit ball of $Y^{\prime}$ into an ordered bounded subset of $(X^{\prime})_{-1}$, then $C$ is $L^{1}$-admissible.
Proof 3.10.
We know from Corollary 2.14(c), that the cone of $(X^{\prime})_{-1}$ is normal. Therefore, we may deduce from Proposition A.1 that there exists an ordered Banach space $\widetilde{X}$ with a unit such that $C^{\prime}$ factorises as
$C^{\prime}:Y^{\prime}\overset{C_{1}}{\longrightarrow}\widetilde{X}\overset{C_{2}}{\longrightarrow}(X^{\prime})_{-1};$
for an operator $C_{1}\in\mathcal{L}(Y^{\prime},\widetilde{X})$ and a positive operator $C_{2}\in\mathcal{L}(\widetilde{X},(X^{\prime})_{-1})$. Taking duals yields
$C^{\prime\prime}:(X_{1})^{\prime\prime}\overset{C_{2}^{\prime}}{\longrightarrow}(\widetilde{X})^{\prime}\overset{C_{1}^{\prime}}{\longrightarrow}Y^{\prime\prime};$
here we have used the equality $(X^{\prime})_{-1}=(X_{1})^{\prime}$. Restricting ourselves to $X_{1}$ and employing the reflexivity of $Y$ gives the factorisation
$C:X_{1}\overset{C_{2}^{\prime}}{\longrightarrow}(\widetilde{X})^{\prime}\overset{C_{1}^{\prime}}{\longrightarrow}Y,$
where $C_{2}^{\prime}$ is positive because $C_{2}$ is.
Finally, since $\widetilde{X}$ has a unit, so the norm on $(\widetilde{X})^{\prime}$ is additive on its positive cone [42, Lemmata 2 and 3]. We have thus shown that all assumptions of Theorem 3.5 are fulfilled and thus $C$ is $L^{1}$-admissible.
3.2 Zero-class $L^{1}$-admissibility
While $L^{1}$-admissible observation operators need not be zero-class $L^{1}$-admissible (see Example 3.16 below), we are interested in situations when this is indeed the case. To this end, we take advantage of the nilpotency of the right-translation semigroup on $L^{1}([0,1],Y)$ to formulate a simple test for $L^{1}$-admissibility to be zero-class.
Proposition 3.11.
Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$, let $Y$ be a Banach space, and let $(R_{t})_{t\geq 0}$ denote the nilpotent right-translation semigroup on $L^{1}([0,1],Y)$.
Suppose that $C\in\mathcal{L}(X_{1},Y)$ is $L^{1}$-admissible. Then $C$ is zero-class $L^{1}$-admissible if and only if $\lim_{t\uparrow 1}\left\lVert R_{t}\Psi_{1}\right\rVert_{L^{1}}=0$, where $\Psi_{(\mathord{\,\cdot\,})}$ is the output map defined in (3.1).
Proof 3.12.
For each $\tau<1$ and $x\in X_{1}$, we have
$\left\lVert R_{1-\tau}\Psi_{1}x\right\rVert_{L^{1}}$ $=\int_{0}^{1}\left\lVert(R_{1-\tau}\Psi_{1}x)(s)\right\rVert_{Y}\;\mathrm{d}s=\int_{1-\tau}^{1}\left\lVert(\Psi_{1}x)(s-1+\tau)\right\rVert_{Y}\;\mathrm{d}s$
$=\int_{1-\tau}^{1}\left\lVert CT(s-1+\tau)x\right\rVert_{Y}\;\mathrm{d}s=\int_{0}^{\tau}\left\lVert CT(s)x\right\rVert_{Y}\;\mathrm{d}s=\left\lVert\Psi_{\tau}x\right\rVert_{L^{1}}.$
The assertion is now immediate.
A subset $S$ of an ordered Banach space $X$ is called almost order-bounded if for each $\epsilon>0$ there exists $x_{\epsilon}\in X_{+}$ such that $\left\lVert(\left\lvert x\right\rvert-x_{\epsilon})_{+}\right\rVert<\epsilon$ for all $x\in S$. Equivalently, if for each $\epsilon>0$ there exists $x_{\epsilon}\in X_{+}$ such that
$S\subseteq[-x_{\epsilon},x_{\epsilon}]+\epsilon B_{X};$
where $B_{X}$ denotes the unit ball of $X$. Moreover, an AL-space is a Banach lattice whose norm is additive on the positive cone (cf. Table 1).
Theorem 3.13.
Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a reflexive Banach space $X$ and let $Y$ be an AL-space. If $C\in\mathcal{L}(X_{1},Y)$ is $L^{1}$-admissible, then it is also zero-class $L^{1}$-admissible.
Proof 3.14.
Let $(R_{t})_{t\geq 0}$ denote the nilpotent right-translation semigroup on the space $L^{1}([0,1],Y)$. Reflexivity of $X$ implies weak compactness of the output operator $\Psi_{1}$. In turn, $\Psi_{1}(B_{X})$ is relatively weakly compact. Next, as $Y$ is an AL-space, so is $L^{1}([0,1],Y)$. Consequently, $\Psi_{1}(B_{X})$ is almost order-bounded by Dunford-Pettis theorem [38, Theorem 2.5.4]. Strong continuity and nilpotency of $(R_{t})_{t\geq 0}$ now implies that it converges to 00 uniformly on $\Psi_{1}(B_{X})$ as $t\uparrow 1$. Consequently, $C$ is zero-class $L^{1}$-admissible by Lemma 3.11.
Note that all of the order assumptions in Theorem 3.13 are posed on $Y$ whereas $X$ is a general Banach space. We don’t know whether the Dunford-Pettis theorem [38, Theorem 2.5.4] can be generalised to ordered Banach spaces whose norm is additive on the positive cone. For this reason, the proof above necessitated that $Y$ be a Banach lattice.
Remark 3.15.
In the proof of the above theorem, reflexivity is only used to obtain weak compactness of $\Psi_{1}$. In other words, for $L^{1}$-admissible observation operators mapping into an AL-space, weak compactness of the output operator implies zero-class $L^{1}$-admissibility.
The following example shows that the reflexivity assumption in Theorem 3.13 cannot be dropped, in general; see also [33, Example 26].
Example 3.16.
Consider the left translation semigroup on $X:=L^{1}[0,1]$ with observation $C:X_{1}\to\mathbb{C}$ given by $f\mapsto f(0)$. The corresponding input operator $\Psi_{\tau}:X_{1}\to L^{1}([0,\tau],Y)$ satisfies $\left\lVert\Psi_{\tau}\right\rVert=1$ for each $\tau>0$. Therefore, $C$ is $L^{1}$-admissible but the admissibility is not zero-class.
Combining Proposition 3.1 with Theorem 3.13 yields the following sufficient condition for zero-class $L^{1}$-admissibility of positive observation operators:
Corollary 3.17.
Suppose that Assumption 3 holds and assume that $X$ is reflexive and $Y$ is an AL-space. If $C\in\mathcal{L}(X_{1},Y)$ is positive, then it is zero-class $L^{1}$-admissible.
Yet another condition for zero-class $L^{1}$-admissibility of observation operators is given in Lemma 5.5.
4 Admissibility of control operators
In this section, we shift our attention to the admissibility considerations of the system
[…]
We make the following assumption on the $C_{0}$-semigroup $(T(t))_{t\geq 0}$ generated by $A$ and the control operator $B\in\mathcal{L}(U,X_{-1})$:
{assumption}
Suppose that $X$ is an ordered Banach space with a generating and normal cone and $(T(t))_{t\geq 0}$ is a positive $C_{0}$-semigroup on $X$. Moreover, $U$ is a Banach space and $B\in\mathcal{L}(U,X_{-1})$.
In what follows, we let $\mathrm{T}([0,\tau],U)$ denote the space of all $U$-valued step functions on $[0,\tau]$, i.e., the functions that can be written as a finite linear combination of characteristic functions of intervals. This is a normed space when equipped with the supremum norm whose closure is called the space of regulated functions – denoted by $\operatorname{Reg}([0,\tau],U)$.
Now, let $\mathrm{Z}$ be a placeholder for $L^{\infty}$ or $\mathrm{C}$ or $\operatorname{Reg}$. Corresponding to the system in (4.1), we define the input map as
(4.2) $\Phi_{\tau}:\mathrm{Z}([0,\tau],U)\to X_{-1},\qquad u\mapsto\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s$
for fixed $\tau>0$. The operator $B$ is called $\mathrm{Z}$-admissible control operator if $\operatorname{Rg}\Phi_{\tau}\subseteq X$ for some (equivalently, all) $\tau>0$. Additionally, if $\limsup_{\tau\downarrow 0}\left\lVert\Phi_{\tau}\right\rVert_{\mathcal{L}(\mathrm{Z}([0,\tau],U),X)}=0$, then we say that $B$ is a zero-class $\mathrm{Z}$-admissible control operator.
Since the space of regulated functions contains the space of continuous functions, so every $\operatorname{Reg}$-admissible control operator is $\mathrm{C}$-admissible. In fact, the converse is also true and seems to be folklore. The proof for the particular case $U=X$ and $B=A_{-1}$ is given in [32, Proposition 2.2] and the proof for the general case follows mutatis mutandis. As we use this observation multiple times in the sequel, we state it explicitly:
Proposition 4.1.
Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a Banach space $X$, let $U$ be a Banach space, and let $B\in\mathcal{L}(U,X_{-1})$. The operator $B$ is $\mathrm{C}$-admissible if and only if $B$ is $\operatorname{Reg}$-admissible.
4.1 Positivity of $B$ in terms of the boundary control operator
As explained in the introduction, the additive input term $Bu$ in (4.1) with an operator $B:U\to X_{-1}$ is often used to encode the boundary control. Since several results in this section use the assumption that $B$ is positive, we will first describe how the positivity of $B$ can be described in terms of the boundary operator. The precise setting is as follows.
Let $X$ and $U$ be ordered Banach spaces and let $\mathfrak{A}:\operatorname{dom}\mathfrak{A}\subset X\to X$ and $\mathfrak{B}:\operatorname{dom}\mathfrak{A}\to U$ be linear operators. Assume that $\mathfrak{B}$ has a bounded right-inverse and that the restriction $A:=\mathfrak{A}|_{\ker\mathfrak{B}}$ generates a positive $C_{0}$-semigroup on $X$. For every $\lambda\in\rho(A)$, one has the following properties (see for instance [24, Lemma 1.2]): the domain $\operatorname{dom}\mathfrak{A}$ is the direct sum of its vector subspaces $\operatorname{dom}(A)=\ker\mathfrak{A}$ and $\ker(\lambda-\mathfrak{A})$ and the restriction $\mathfrak{B}|_{\ker(\lambda-\mathfrak{A})}$ is a bijection from $\ker(\lambda-\mathfrak{A})$ to $U$. We denote its inverse by $B_{\lambda}:U\to\ker(\lambda-\mathfrak{A})\subseteq\operatorname{dom}\mathfrak{A}\subseteq X$. The boundary control problem
$\dot{x}(t)={}$ $\mathfrak{A}x(t),\quad t>0,\quad x(0)=0,$
$\mathfrak{B}x(t)={}$ $u(t)$
with control $u(t)\in U$ can, under appropriate assumptions, be reformulated as the system (4.1) if one chooses $B:=(\lambda-A_{-1})B_{\lambda}:U\to X_{-1}$, where $A_{-1}$ is the generator of the extrapolated $C_{0}$-semigroup on $X_{-1}$; see [48, Proposition 2.8]. One can check that the operator $B$ does not depend on the choice of $\lambda$ (see for instance [24, Formula (1.16) in Lemma 1.3]). We now show how positivity of $B$ can be characterised in terms of the operators $B_{\lambda}$.
Proposition 4.2.
In the setting described before the proposition, the following assertions are equivalent:
[…]
    The operator $B:U\to X_{-1}$ is positive.
[…]
    The operator $B_{\lambda}:U\to X$ is positive for all sufficiently large $\lambda\in\mathbb{R}$.
[…]
    The operator $B_{\lambda}:U\to X$ is positive for all $\lambda>s(A)$.
Proof 4.3.
“(i) $\Rightarrow$ (iii)”: Assume that (i) holds and let $\lambda>s(A)$. Then $R(\lambda,A_{-1})$ is a positive operator from $X_{-1}$ to $X$ according to Proposition 2.12(c). Hence, $B_{\lambda}=R(\lambda,A_{-1})B$ is positive.
“(iii) $\Rightarrow$ (ii)”: This implication is obvious.
“(ii) $\Rightarrow$ (i)”: Assume (ii) and let $u\in U_{+}$. Then one has
$Bu=\lim_{\lambda\to\infty}\lambda R(\lambda,A_{-1})Bu=\lim_{\lambda\to\infty}\lambda B_{\lambda}u\in X_{-1,+},$
where both limits are norm limits in $X_{-1}$.
We point out that the equivalence of (i) and (iii) in Proposition 4.2 above was also noted in [19, Lemma 2.1] and on [20, Page 16].
In a similar vein, if one already knows that a boundary control system is admissible, then the positivity of the operators $B_{\lambda}$ characterises whether positive inputs lead to positive trajectories [16, Proposition 4.3]. The condition that all $B_{\lambda}$ be positive also occurs in assumption (ii) of [19, Theorem 4.3]. It is instructive to discuss what Proposition 4.2 says for the Laplace operator. The following example demonstrates that the positivity condition on $B_{\lambda}$ can be interpreted as the maximum principle for the Laplace operator in this case.
Example 4.4.
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded smooth domain (say, for the sake of simplicity, with $C^{\infty}$-boundary), set $X:=L^{p}(\Omega)$ for some $p\in(1,\infty)$ and let $\mathfrak{A}$ denote the Laplace operator with domain $W^{2,p}(\Omega)$. Let $U\subseteq L^{p}(\partial\Omega)$ denote the image of the trace operator on $W^{2,p}(\Omega)$, which we endow with the order inherited from $L^{p}(\partial\Omega)$, and let $\mathfrak{B}:W^{2,p}(\Omega)\to U$ be the trace operator.
Then $A$ is the Dirichlet Laplace operator on $L^{p}(\Omega)$. For $\lambda>s(A)$ and $u\in U$, the function $v:=B_{\lambda}u\in W^{2,p}(\Omega)$ solves the boundary value problem
$(\lambda-\Delta)v$ $=0,$
$v|_{\partial\Omega}$ $=u.$
So by the maximum principle, the operator $B_{\lambda}$ is positive. Hence, Proposition 4.2 yields that $B$ is positive.
4.2 Input admissibility for positive semigroups
If $U$ and $X$ in Assumption 4 are reflexive and if $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$, then the same assumption holds for the double dual $B^{\prime\prime}\in\mathcal{L}(U^{\prime\prime},(X^{\prime\prime})_{-1})$. Due to the reflexivity of $U$, we obtain that $B^{\prime}$ is an $L^{1}$-admissible observation operator by Corollary 3.9. Therefore, reflexivity of $X$ allows us to appeal to the Weiss duality theorem [55, Theorem 6.9(ii)] to obtain that $B$ is $L^{\infty}$-admissible. Without the reflexivity assumptions on $U$ and $X$, we are able to show that $L^{\infty}$-admissibility is always zero-class.
Theorem 4.5.
Suppose that Assumption 4 holds and that $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. If $B$ is $L^{\infty}$-admissible, then it is zero-class $L^{\infty}$-admissible.
In particular, $B$ is always zero-class $L^{\infty}$-admissible if both $X$ and $U$ are reflexive.
Proof 4.6.
By assumption, there exist $a,b\in X_{-1}$ such that $B(B_{U})\subseteq[a,b]$. Without loss of generality, assume that $\omega_{0}(A)<0$ and that the norm on $X_{-1}$ is given by $\left\lVert x\right\rVert_{-1}=\left\lVert(A_{-1})^{-1}x\right\rVert_{X}$ for all $x\in X_{-1}$. Suppose that $B$ is $L^{\infty}$-admissible. Then the corresponding input operator $\Phi_{\tau}:L^{\infty}([0,\tau],U)\to X_{-1}$ maps into $X$ and for each $u\in L^{\infty}([0,\tau],U)$ taking values in $B_{U}$, we get
$\left\lVert\Phi_{\tau}u\right\rVert_{X}\leq\left\lVert\int_{0}^{\tau}T_{-1}(s)b\;\mathrm{d}s\right\rVert_{X}=\left\lVert A_{-1}\int_{0}^{\tau}T_{-1}(s)b\;\mathrm{d}s\right\rVert_{-1}=\left\lVert T_{-1}(\tau)b-b\right\rVert_{-1}\to 0$
as $\tau\to 0$. In other words, $B$ is zero-class $L^{\infty}$-admissible. The second assertion can now be inferred from the discussion preceding the theorem.
Remark 4.7.
With the same arguments as above, it can be shown that the first assertion of Theorem 4.5 remains true if $L^{\infty}$ is replaced with $\mathrm{C}$.
Dropping the reflexivity assumptions in the second part of Theorem 4.5 we are able to at least show zero-class $\mathrm{C}$-admissibility:
Theorem 4.8.
Suppose that Assumption 4 holds and that $B\in\mathcal{L}(U,X_{-1})$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. Then $B$ is zero-class $C$-admissible.
Proof 4.9.
For each $u\in\mathrm{T}([0,\tau],U)$ – the space of all $U$-valued step functions on $[0,\tau]$ – we know that
$\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\in X$
from [10, Lemma 4.3(ii)]; note that the proof of [10, Lemma 4.3(ii)] doesn’t use any positivity properties. The positivity of the semigroup and the assumption on $B$ guarantees that there exists $b\in X_{-1}$ such that if $u\in\mathrm{T}([0,\tau],U)$ with $\left\lVert u\right\rVert_{\infty}\leq 1$, then
$\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert\leq\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)b\;\mathrm{d}s\right\rVert=:K_{\tau}$
As a result,
$\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert\leq K_{\tau}\left\lVert u\right\rVert_{\infty}$
for all step functions $u\in\mathrm{T}([0,\tau],U)$. Hence, $B$ is $\mathrm{C}$-admissible (see, [56, Remark 4.1.7] and Proposition 4.1). The zero-class $\mathrm{C}$-admissibility now follows from Remark 4.7.
In particular, the following generalisation of [56, Corollary 4.2.10] follows by combining Theorem 4.8 and Proposition A.3(a).
Corollary 4.10.
Suppose that Assumption 4 holds and let $U$ be an ordered Banach space with a unit. Then every positive $B\in\mathcal{L}(U,X_{-1})$ is zero-class $\mathrm{C}$-admissible.
Another situation where $B$ satisfies the assumptions of Theorem 4.8 is given in Proposition A.3(b). However, such a condition even yields zero-class $L^{\infty}$-admissibility as we see in Corollary 4.16.
Remark 4.11.
The assumption that $U$ has a unit cannot be dropped in Corollary 4.10. Indeed, on the Banach lattice $c_{0}$, there is an example [32, Example 2.3] – due to Kato – of a positive semigroup whose generator is positive, $\mathrm{C}$-admissible, yet it is not zero-class $\mathrm{C}$-admissible.
Recall from Table 1 that an AM-space is a Banach lattice whose closed unit ball is upwards directed. In particular, $c_{0}$ is an AM-space. It was shown in [56, Theorem 4.1.19] that if $U$ is an AM-space, then every positive $B\in\mathcal{L}(U,X_{-1})$ is at least $\mathrm{C}$-admissible. We generalise this below:
Theorem 4.12.
Suppose that Assumption 4 holds and let $U$ be an ordered Banach space whose closed unit ball is upwards directed. Then every positive operator $B\in\mathcal{L}(U,X_{-1})$ is $\mathrm{C}$-admissible.
Proof 4.13.
Without loss of generality, we assume that $\omega_{0}(A)<0$. Let $B\in\mathcal{L}(U,X_{-1})$ be positive and let $\mathrm{T}([0,\tau],U)$ denote the space of all $U$-valued step functions on $[0,\tau]$. Let $\Phi_{\tau}$ be the input operator in (4.2) with $Z=\operatorname{Reg}$. As in the proof of Theorem 4.8, $\Phi_{\tau}$ maps $\mathrm{T}([0,\tau],U)$ into $X$.
Let $u\in\mathrm{T}([0,\tau],U)$. Then we can write $u=\sum_{k=1}^{n}u_{k}\chi_{I_{k}}$ for $u_{k}\in U$ and disjoint intervals $I_{k}\subseteq[0,\tau]$. Since $B_{U}$ is upwards directed, the norm on $X^{\prime}$ is additive on the positive cone $(X^{\prime})_{+}$ by [42, Lemma 3]. Whence, according to [37, Theorem 1.3(2)], there exist $\alpha>1$ – independent of $u$ – and $0\leq w\in U$ such that $u_{k}\leq w$ for each $k$ and $\left\lVert w\right\rVert\leq\alpha\left\lVert u\right\rVert_{\infty}$. In particular, $u(s)\leq w$ for all $s\in[0,\tau]$.
Positivity of the semigroup, the operator $B$, and the element $w\in U$ now allow us to estimate $\left\lVert\int_{0}^{\tau}T_{-1}(\tau-s)Bu(s)\;\mathrm{d}s\right\rVert$ from above by
$\left\lVert\int_{0}^{\infty}T_{-1}(\tau-s)Bw\;\mathrm{d}s\right\rVert\leq\left\lVert A_{-1}^{-1}B\right\rVert\left\lVert w\right\rVert\leq\alpha\left\lVert A_{-1}^{-1}B\right\rVert\left\lVert u\right\rVert_{\infty}.$
It follows that $B$ is $\operatorname{Reg}$-admissible due to [56, Remark 4.1.7]. The $\mathrm{C}$-admissibility is therefore true due to Proposition 4.1.
Since the closed unit ball is upwards directed if and only if the dual norm is additive on the positive cone [52, Proposition 7 and Theorem 8], Proposition 3.1 and Theorem 4.12 are dual to each other.
Theorem 4.14.
Suppose that Assumption 4 holds and assume that $X_{+}$ is a face of $X^{\prime\prime}_{+}$. If $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$, then $B$ is zero-class $L^{\infty}$-admissible.
Proof 4.15.
By Theorem 4.5, it suffices to show that $B$ is $L^{\infty}$-admissible. By assumption, there exists $a,b\in X_{-1}$ such that $B(B_{U})\subseteq[a,b]$. First of all, note that $\int_{0}^{\tau}T_{-1}(\tau-s)a\;\mathrm{d}s,\int_{0}^{\tau}T_{-1}(\tau-s)b\;\mathrm{d}s\in\operatorname{dom}A_{-1}=X$. Now for each $u\in L^{\infty}([0,\tau],U)$ with $\left\lVert u\right\rVert_{\infty}\leq 1$, we have
$\int_{0}^{\tau}T_{-1}(\tau-s)a\;\mathrm{d}s\leq\Phi_{\tau}u\leq\int_{0}^{\tau}T_{-1}(\tau-s)b\;\mathrm{d}s.$
Because $X_{+}$ is a face of $X^{\prime\prime}_{+}$, also $X_{+}$ is a face of $X_{-1,+}$ by Corollary 2.15. It follows that $\Phi_{\tau}u\in X$. A rescaling argument yields that $\Phi_{\tau}u\in X$ for all $u\in L^{\infty}([0,\tau],U)$.
We give some sufficient conditions in Proposition A.3 under which a positive control operator automatically satisfies the assumptions of Theorem 4.14. In particular, zero-class $L^{\infty}$-admissibility is automatic in these situations:
Corollary 4.16.
Suppose that Assumption 4 holds and let $U$ be an ordered Banach space and let $B\in\mathcal{L}(U,X_{-1})$ be positive.
The following assumptions are sufficient for $B$ to be zero-class $L^{\infty}$-admissible.
[…]
    The space $U$ has a unit and $X_{+}$ is a face of $X^{\prime\prime}_{+}$,
[…]
    The unit ball of $U$ is upwards directed and every increasing norm-bounded net in $X_{+}$ is norm convergent.
Proof 4.17.
In each case, $X_{+}$ is a face of $X^{\prime\prime}_{+}$, so by Theorem 4.14, we only need to show that $B$ maps the unit ball of $U$ into an order-bounded subset of $X_{-1}$. For (a), this is a direct consequence of Proposition A.3(a).
(b) Due to Corollary 2.14(a), every increasing norm-bounded net in $X_{-1,+}$ is norm-convergent. The claim now follows by Proposition A.3(b).
Here too, the example [32, Example 2.3] by Kato mentioned in Remark 4.11 shows that the assumption that $U$ has a unit cannot be dropped in Corollary 4.16(a).
4.3 Input admissibility for general semigroups
In this section, we leave the setting of Assumption 4 and instead impose order assumptions on the input space. This allows us to give conditions under which admissibility of the control operator automatically gives zero-class $L^{\infty}$-admissibility. First, we obtain the following analogue to Theorem 3.13:
Theorem 4.18.
Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a reflexive Banach space $X$ and let $U$ be an AM-space. If $B\in\mathcal{L}(U,X_{-1})$ is $L^{\infty}$-admissible, then it is also zero-class $L^{\infty}$-admissible.
Proof 4.19.
Let $B$ be $L^{\infty}$-admissible. Then by [55, Theorem 6.9(iii)], we know that $B^{\prime}$ is $L^{1}$-admissible. Since $U$ is an AM-space, we have that $U^{\prime}$ is an AL-space. Moreover, as $X$ is reflexive, so is $X^{\prime}$. Thus, Theorem 3.13 implies that $B^{\prime}$ is also zero-class $L^{1}$-admissible. Keeping in mind that $X$ is reflexive, the result follows by [55, Theorem 6.9(ii)].
In fact, if $U$ is finite-dimensional, then even $\mathrm{C}$-admissibility of the control operator implies that it is zero-class $L^{\infty}$-admissibility:
Theorem 4.20.
Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a reflexive Banach space $X$ and let $U$ be finite-dimensional. If $B\in\mathcal{L}(U,X_{-1})$ is $\mathrm{C}$-admissible, then it is also zero-class $L^{\infty}$-admissible.
For the proof of the theorem, we need the following lemma in the spirit of Weiss duality results [55, Theorem 6.9]:
Lemma 4.21.
Let $(T(t))_{t\geq 0}$ be a $C_{0}$-semigroup on a reflexive Banach space $X$ and let $U$ be finite-dimensional. If $B\in\mathcal{L}(U,X_{-1})$ is a $\mathrm{C}$-admissible control operator, then $B^{\prime}\in\mathcal{L}((X^{\prime})_{1},U)$ is an $L^{1}$-admissible observation operator.
Proof 4.22.
Let $\tau>0$. Since $B$ is $\mathrm{C}$-admissible, there exists $K_{B,\tau}>0$ such that
$\left\lVert\int_{0}^{\tau}T_{-1}(t)Bu(t)\;\mathrm{d}t\right\rVert\leq K_{B,\tau}\left\lVert u\right\rVert_{\infty}$
for all $u\in\mathrm{C}([0,\tau],U)$. Finite-dimensionality of the input space allows us to compute
$\left\lVert B^{\prime}T^{\prime}(\mathord{\,\cdot\,})x^{\prime}\right\rVert_{1}$ $=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\int_{0}^{\tau}\left\langle B^{\prime}T^{\prime}(t)x^{\prime}\,,\,u(t)\right\rangle\right\rvert\;\mathrm{d}t$
$=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\int_{0}^{\tau}\left\langle B^{\prime}(T_{-1})^{\prime}(t)x^{\prime}\,,\,u(t)\right\rangle\right\rvert\;\mathrm{d}t$
$=\sup_{u\in\mathrm{C}([0,\tau],U),\left\lVert u\right\rVert_{\infty}\leq 1}\left\lvert\left\langle x^{\prime}\,,\,\int_{0}^{\tau}T_{-1}(t)Bu(t)\;\mathrm{d}t\right\rangle\right\rvert$
$\leq K_{B,\tau}\left\lVert x^{\prime}\right\rVert$
for all $x^{\prime}\in\operatorname{dom}A^{\prime}$. This proves that $B^{\prime}$ is $L^{1}$-admissible.
Proof 4.23 (Proof of Theorem 4.20).
Let $B$ be $\mathrm{C}$-admissible. Since $X$ is reflexive and $U$ is finite-dimensional, we have that $B^{\prime}$ is $L^{1}$-admissible by Lemma 4.21. It follows from Theorem 3.13 that $B^{\prime}$ is also zero-class $L^{1}$-admissible. Once again, using the reflexivity of $X$, it follows that $B$ is zero-class $L^{\infty}$-admissible by the Weiss duality result [55, Theorem 6.9(ii)].
4.4 $L^{r}$-input admissibility of positive semigroups
We close this section by moving a bit away from the limit-cases of $L^{\infty}$- and $C$-admissibility and instead considering $L^{r}$-admissibility for $r<\infty$. If one has a sufficiently strong estimate on the resolvent one can even get $L^{1}$-admissibility of positive input operators [19, Theorems 2.1]; compare however Appendix B at the end of this paper. In the following corollary, we work with an intrinsic ultracontractivity type assumption on the semigroup instead which was shown in Theorem 2.18(b) to imply that $\operatorname{span}X_{-1,+}$ is contained in a negative Favard space.
Corollary 4.24.
Let $(T(t))_{t\geq 0}$ be a positive and analytic semigroup on a reflexive ordered Banach space $X$. Let $\varphi\in X^{\prime}$ be strictly positive, assume that $\operatorname{dom}A^{\prime}\cap(X^{\prime})_{\varphi}$ is majorizing in $(X^{\prime})_{\varphi}$, that $T(t)^{\prime}X^{\prime}\subseteq(X^{\prime})_{\varphi}$ for each $t>0$ and that the ultracontractivity type estimate $\left\lVert T(t)^{\prime}\right\rVert_{X^{\prime}\to(X^{\prime})_{\varphi}}\leq\widetilde{c}t^{-\alpha}$ is satisfied for constants $\widetilde{c}\geq 0$ and $\alpha\in[0,1)$ and for all $t\in(0,1]$.
If $U$ is an ordered Banach space and $B\in\mathcal{L}(U,X_{-1})$ is positive (or, more generally, the difference of two positive operators), then $B$ is $L^{r}$-admissible for every $r>\frac{1}{1-\alpha}$.
Proof 4.25.
By Theorem 2.18(b), $B$ maps into the Favard space $F_{-\alpha}$, whence, by the closed graph theorem, $B\in\mathcal{L}(U,F_{-\alpha})$. This readily implies that $B$ is $L^{r}$-admissible for $r>\frac{1}{1-\alpha}$, see e.g. [45, Lemma 2.1 and Remark 2.3] or [35, Proposition 19].
Example 4.26.
Let $\Omega\subseteq\mathbb{R}^{d}$ be a bounded domain with Lipschitz boundary, let $p,q\in(1,\infty)$ be Hölder conjugates and let $A$ be the elliptic operator with Neumann boundary conditions on $L^{p}(\Omega)$ from Example 2.20. If $\alpha:=\frac{d}{2q}<1$, then it follows from assertion (a) in Example 2.20 that $\operatorname{span}L^{p}(\Omega)_{-1,+}$ is contained in the Favard space $F_{-\alpha}$.
Now let $U$ be an ordered Banach space and let $B\in\mathcal{L}(U,X_{-1})$ be positive or, more generally, the difference of two positive operators. Than $B$ is $L^{r}$-admissible for every $r\in(\frac{1}{1-\alpha},\infty]$. If $s\in[1,\infty)$ denote the Hölder conjugate of $r$, this can be reformulated by saying that $B$ is $L^{r}$-admissible if $s<\frac{2q}{d}=\frac{1}{\alpha}$.
We note that the derived range of $r$ may not be optimal, partly due to its dependence on the dimension $d$. Indeed, it is known that the operator $B$ given by Neumann boundary control is $L^{r}$-admissible for $r>4/3$, if the boundary of $\Omega$ is $C^{2}$, see e.g. [34] or [48, Example 2.14] and even $r\geq 4/3$ when the boundary is $C^{\infty}$ [27, Proposition 2.4] and [45, Remark 2.8].
5 Applications to perturbation results
The consequences of admissibility for perturbation results are well-known. Indeed, if $(T(t))_{t\geq 0}$ is a $C_{0}$-semigroup on a Banach space and $B\in\mathcal{L}(X,X_{-1})$ is zero-class $\mathrm{C}$-admissible, then the restriction of $A_{-1}+B$ to $X$ generates a $C_{0}$-semigroup on $X$ [17, Corollary III.3.3]. Likewise, zero-class $L^{1}$-admissibility of the observation operator also yields a perturbation result [17, Corollary III.3.16]. Using results obtained in the prequel, we are thus able to obtain perturbation results for positive $C_{0}$-semigroups.
Corollary 5.1.
Let $X$ be an ordered Banach space with a generating and normal cone and let $A$ be the generator of a positive $C_{0}$-semigroup on $X$.
If $B\in\mathcal{L}(X,X_{-1})$ maps the unit ball of $X$ to an order-bounded set in $X_{-1}$, then the part of $A_{-1}+B$ in $X$ generates a $C_{0}$-semigroup on $X$. If $B$ is positive, so is the perturbed semigroup.
Proof 5.2.
The assumptions imply, according to Theorem 4.8, that $B$ is zero-class $\mathrm{C}$-admissible. Hence, the claim follows from the perturbation result in [17, Corollary III.3.3].
As a consequence of Corollary 5.1, one can at once see that the operator $A_{-1}+B$ in [10, Example 5.1] generates a $C_{0}$-semigroup without having to check any spectral conditions and one can also allow multiplication operator with function in $L^{1}$ as perturbations:
Example 5.3.
On the space $X=\{f\in C[0,1]:f(1)=0\}$ the operator $A$ given by
$\operatorname{dom}A:=\{f\in C^{1}[0,1]:f(1)=f^{\prime}(1)=0\},\qquad Af:=f^{\prime}$
generates a positive $C_{0}$-semigroup [17, Example II.3.19(i)] that corresponds to the partial differential equation
$u_{t}(t,x)$ $=u_{x}(t,x)$ for $x\in[0,1],t\geq 0,$
$u(0,x)$ $=u_{0}(x)$ for $x\in[0,1],$
$u(t,1)$ $=0$ for $t\geq 0.$
[…]
Either by directly dualizing Corollary 5.1 or by using the results from Section 4 and then dualizing one also gets the following perturbation result:
Corollary 5.4.
Let $X$ be a reflexive ordered Banach space with a generating and normal cone and let $A$ be the generator of a positive $C_{0}$-semigroup on $X$.
If $C\in\mathcal{L}(X_{1},X)$ is such that $C^{\prime}$ maps the unit ball of $X^{\prime}$ to an order-bounded set in $(X^{\prime})_{-1}$, then $A+C$ generates a $C_{0}$-semigroup on $X$. If $C$ is positive, so is the perturbed semigroup.
The perturbation assertions of Corollary 5.4 hold once we establish that $C$ is zero-class $L^{1}$-admissible [17, Corollary III.3.16]. Let us prove this separately in the spirit of Section 3.2:
Lemma 5.5.
Let $X$ be a reflexive ordered Banach space with a generating and normal cone, let $(T(t))_{t\geq 0}$ be a positive $C_{0}$-semigroup on $X$, and let $Y$ be a Banach space. If $C\in\mathcal{L}(X_{1},Y)$ is such that $C^{\prime}$ maps the unit ball of $Y^{\prime}$ to an order-bounded set in $(X^{\prime})_{-1}$, then $C$ is zero-class $L^{1}$-admissible.
Proof 5.6.
Since $X$ is reflexive, $X_{+}$ is a face of $X^{\prime\prime}_{+}$. Therefore, $C^{\prime}\in\mathcal{L}(Y^{\prime},X^{\prime}_{-1})$ is zero-class $L^{\infty}$-admissible by Theorem 4.14. Zero-class $L^{1}$-admissibility of $C$ is now a consequence of Weiss duality result [55, Theorem 6.9(a)].
Note that this theorem implies in particular the following result: every positive finite rank perturbation $C:X_{1}\to X$ (and more generally, every finite rank perturbation that can be written as the difference of two positive operators $X_{1}\to X$) of a positive $C_{0}$-semigroup on a reflexive Banach lattice again generates a $C_{0}$-semigroup. In fact, this is known to be true even if $X$ is not reflexive and goes back to Arendt and Rhandi [8, Corollary 2.4]. Note that this is not true, in general, for finite rank perturbations that do not satisfy any positivity assumption: Desch and Schappacher proved in [15, Theorem 2] that a $C_{0}$-semigroup on a Banach space $X$ is automatically analytic if every rank-$1$ perturbation $C:X_{1}\to X$ still generates a $C_{0}$-semigroup.
Appendix A order-bounded images of the unit ball
Let $U$ be a Banach space and $Z$ be an ordered Banach space. In Section 4, the property that an operator $T\in\mathcal{L}(U,Z)$ maps the unit ball of $U$ into an order-bounded subset of $Z$ played an important role in obtaining sufficient conditions for admissibility. In this appendix, we give a characterisation and, for positive $T$, a number of sufficient conditions for this property. We note in passing that if $Z$ is even a Banach lattice, the slightly weaker property that $T$ maps the unit ball of $U$ into an order-bounded subset of $Z^{\prime\prime}$ is called majorizing [47, Proposition IV.3.4].
Before stating our first result, we clarify that by a null sequence in a Banach space, we mean a sequence converging to 00. Additionally, we make the following simple observation: let $S$ be an order-bounded subset of an ordered Banach space $X$ such that $0\in S$. Then there exists $x,z\in X_{+}$ such that $S\subseteq[-x,z]$. Taking $e=x+z\in X_{+}$, we get that $S$ is contained in $[-e,e]$.
Proposition A.1.
Let $U$ be a Banach space and let $Z$ be an ordered Banach space with a normal cone. The following conditions are equivalent for an operator $T\in\mathcal{L}(U,Z)$.
[…]
    The operator $T$ maps the unit ball of $U$ into an order-bounded subset of $Z$.
[…]
    There exists an ordered Banach space $\widetilde{X}$ with a unit such that $T$ factorises as
    $T:U\overset{T_{1}}{\longrightarrow}\widetilde{X}\overset{T_{2}}{\longrightarrow}Z$
    for an operator $T_{1}\in\mathcal{L}(U,\widetilde{X})$ and a positive operator $T_{2}\in\mathcal{L}(\widetilde{X},Z)$.
If $Z$ is, in addition, a KB-space, then (i) and (ii) are also equivalent to each of the following conditions:
  1. (iii)
    There exists $c>0$ such that for every null sequence $(u_{n})$ in the open unit ball of $U$, there exists $z\in Z_{+}$ such that $\left\lVert z\right\rVert\leq c$ and $\left\lvert Tu_{n}\right\rvert\leq z$ for all $n\in\mathbb{N}$.
  2. (iv)
    There exists $c>0$ such that $\left\lVert\sup_{u\in F}\left\lvert Tu\right\rvert\right\rVert\leq c\sup_{u\in F}\left\lVert u\right\rVert$ for every finite subset $F\subseteq U$.
Proof A.2.
First, assume that (ii) holds and let $u\in B_{U}$. Since $\widetilde{X}$ has a unit, there exists – according to [22, Proposition 2.11] – an element $e\in\widetilde{X}$ and $\epsilon>0$ such that the implication
$\left\lVert x\right\rVert\leq\epsilon\Rightarrow x\leq e$
holds for all $x\in\widetilde{X}$. In particular, taking $\lambda=\epsilon^{-1}\left\lVert T_{1}\right\rVert$, it follows that $T_{1}u\in[-\lambda e,\lambda e]$; where we have used $\left\lVert u\right\rVert\leq 1$. Employing the positivity of $T_{2}$, it follows that $T=T_{2}T_{1}$ maps the unit ball of $U$ into an order-bounded subset of $Z$.
“(i) $\Rightarrow$ (ii)”: The observation about ordered bounded subsets made prior to the proposition guarantees the existence of $e\in Z_{+}$ such that $T(B_{U})\subseteq[-e,e]$. On the other hand, because the cone of $Z$ is normal, $\widetilde{X}:=Z_{e}$ is an ordered Banach space with a unit. The above argument allows us to infer that $\operatorname{Rg}T\subseteq\widetilde{X}$. Whence, $T$ factors through $\widetilde{X}$ via the canonical embedding $\widetilde{X}\hookrightarrow Z$ (which is positive).
Finally, suppose that $Z$ is a KB-space, then $Z_{+}$ is a face of $(Z^{\prime\prime})_{+}$. The equivalence of (i)-(iv) now follows by [47, Proposition IV.3.4].
Proposition A.3.
Let $U$ and $Z$ be ordered Banach spaces. A positive operator $T\in\mathcal{L}(U,Z)$ maps the unit ball of $U$ into an order-bounded subset of $Z$ in each of the following cases:
[…]
    The space $U$ has a unit.
[…]
    The closed unit ball of $U$ is upwards directed and every increasing norm-bounded net in $Z_{+}$ is norm-convergent.
Proof A.4.
If $U$ has a unit, then due to the positivity of $T$, the assumptions in condition (ii) of Proposition A.1 are fulfilled. Hence, $T$ maps the unit ball of $U$ into an order-bounded subset of $Z$; observe that the proof of (ii) $\Rightarrow$ (i) in Proposition A.1 did not need the cone of $Z$ to be normal.
(b) Let $S=B_{U}\cap U_{+}$. Since $B_{U}$ is upwards directed, for each $u_{1},u_{2}\in S$, there exists $u\in S$ such that $u_{1},u_{2}\leq u$. On the other hand, the assumption on $Z$ implies that the net $(Tu)_{u\in S}$ is norm convergent in $Z$; here we’ve used the positivity of $T$. It follows that $T$ maps the unit ball of $U$ into an order-bounded subset of $Z$.
Appendix B Lower estimate for operators and a characterisation of AL-spaces
Let $E$ be a Banach lattice and let $A:\operatorname{dom}A\subseteq E\to E$ be a resolvent positive operator, i.e., all sufficiently large real numbers $\lambda$ are in the resolvent set of $A$ and satisfy $R(\lambda,A)\geq 0$. In particular, this is satisfied if $A$ generates a positive $C_{0}$-semigroup on $E$. Consider the following condition:
(B.1) There exist numbers $\begin{split}\text{There exist numbers }\lambda>&s(A)\text{ and }c>0\text{ such that }A\\ \text{ satisfies }\left\lVert R(\lambda,A)f\right\rVert&\geq c\left\lVert f\right\rVert\text{ for every }f\in E_{+}.\end{split}$
This condition occurs in a recent paper on infinite-dimensional positive systems [19, Theorems 2.1] where it is shown to have rather strong consequences for admissibility of control operators. Earlier, the condition (B.1) was used to show generation results [5, Theorem 2.5].
The condition (B.1) is, of course, satisfied in the trivial case where $A$ is bounded. More interestingly, if $E$ is an $L^{1}$-space and the semigroup generated by $A$ is stochastic – i.e., positive and norm preserving on $E_{+}$ – then it is also easy to see that (B.1) is satisfied. However, it is not clear at all how to find an unbounded operator on an $L^{p}$-space for $p>1$ that satisfies (B.1). In this appendix we justify this theoretically: in the important special case where $A$ has compact resolvent, the condition (B.1) implies that $E$ is – up to an equivalent renorming – an $L^{1}$-space (Corollary B.3).
Let us recall the following terminology again: An AL-space is a Banach lattice whose norm is additive on the positive cone. Every $L^{1}$-space is an AL-space and conversely, every AL-space can be shown to be isomorphic (as a Banach lattice) to an $L^{1}$-space (over a possible non-$\sigma$-finite measure space). Compare also Table 1 in the introduction. We first show a characterisation of AL-spaces, up to equivalent norms. For related geometric results, see [36, Proposition 3.11] and [2, Section 2].
Theorem B.1.
For each Banach lattice $E$, the following are equivalent:
[…]
    There exists an equivalent norm on $E$ which turns $E$ into an AL-space.
[…]
    No net in the positive unit sphere of $E$ converges weakly to 00.
Proof B.2.
“(i) $\Rightarrow$ (ii)”: This can easily seen by testing against the norm functional on an AL-space.
“(ii) $\Rightarrow$ (i)”: Let (ii) be true. Let $\mathcal{F}^{\prime}$ denote the set of all non-empty finite subsets of the positive unit sphere in $E^{\prime}$. Assume for a moment that the following property ()(*)( ∗ ) holds: For each number $\varepsilon>0$ and each $F^{\prime}\in\mathcal{F}^{\prime}$, there exists a vector $x_{\varepsilon,F^{\prime}}$ in the positive unit sphere of $E$ such that $\langle x^{\prime},x_{\varepsilon,F^{\prime}}\rangle<\varepsilon$ for all $x^{\prime}\in F^{\prime}$. The relation $\preceq$ on $(0,\infty)\times\mathcal{F}^{\prime}$ that is given by $(\varepsilon_{1},F_{1}^{\prime})\preceq(\varepsilon_{2},F_{2}^{\prime})$ if and only if $\varepsilon_{1}\geq\varepsilon_{2}$ and $F_{1}^{\prime}\subseteq F_{2}^{\prime}$, turns $(0,\infty)\times\mathcal{F}^{\prime}$ into a directed set, and the net $(x_{\varepsilon,F^{\prime}})_{(\varepsilon,F^{\prime})\in(0,\infty)\times\mathcal{F}^{\prime}}$ is contained in the positive unit sphere of $E$ and converges weakly to 00. This contradicts (ii), so we conclude that the property ()(*)( ∗ ) does not hold.
Hence, there exists a number $\varepsilon>0$ and a set $F^{\prime}\in\mathcal{F}^{\prime}$ such that for all $x$ in the positive unit sphere of $E$ one has $\langle x^{\prime},x\rangle\geq\varepsilon$ for at least one element $x^{\prime}\in F^{\prime}$. So if we set $z^{\prime}:=\sum_{x^{\prime}\in F^{\prime}}x^{\prime}$, then we obtain $\langle z^{\prime},x\rangle\geq\varepsilon\left\lVert x\right\rVert$ for each $x\in E_{+}$. Hence, $x\mapsto\langle z^{\prime},\left\lvert x\right\rvert\rangle$ is an equivalent lattice norm on $E$, which is clearly additive on the positive cone.
Corollary B.3.
Let $E$ be a Banach lattice and let $T:E\to E$ be a compact linear operator. If there exists a number $c>0$ such that $\left\lVert Tx\right\rVert\geq c\left\lVert x\right\rVert$ for all $x\in E_{+}$, then there exists an equivalent norm on $E$ which turns $E$ into an AL-space.
Proof B.4.
Assume that there is no equivalent norm on $E$ which turns $E$ into an AL-space. Then there exists, by Theorem B.1, a net $(x_{j})$ in the positive unit sphere of $E$ which converges weakly to 00. Since $T$ is compact, it turns weakly convergent nets into norm convergent ones, so $(Tx_{j})$ converges to 00 in norm. However, one has $\left\lVert Tx_{j}\right\rVert\geq c$ for each index $j$ by assumption; this is a contradiction.
We find it instructive to state the following special case of Corollary B.3.
Corollary B.5.
Let $E$ be a Banach lattice and let $A:\operatorname{dom}A\subseteq E\to E$ be a resolvent positive linear operator. If $A$ has compact resolvent and satisfies (B.1), then there exists an equivalent norm on $E$ which turns $E$ into an AL-space.
Acknowledgements
The first author was funded by the Deutsche Forschungsgemeinschaft (DFG, German Research Foundation) – 523942381. The article was initiated during a pleasant visit of the first author to the third author at the Tampere University, Finland in autumn 2023. The first author is indebted to COST Action 18232 for financial support for this visit. The third author was funded by the Research Council of Finland grant number 349002.